### v1

# Synomaly Noise and Multi-Stage Diffusion: A Novel Approach for Unsupervised Anomaly Detection in Ultrasound Imaging

Abstract
Ultrasound (US) imaging is widely used in routine clinical practice due to its advantages of being radiation-free, cost-effective, and portable. However, the low reproducibility and quality of US images, combined with the scarcity of expert-level annotation, make the training of fully supervised segmentation models challenging. To address these issues, we propose a novel unsupervised anomaly detection framework based on a diffusion model that incorporates a synthetic anomaly (Synomaly) noise function and a multi-stage diffusion process. Synomaly noise introduces synthetic anomalies into healthy images during training, allowing the model to effectively learn anomaly removal. The multi-stage diffusion process is introduced to progressively denoise images, preserving fine details while improving the quality of anomaly-free reconstructions. The generated high-fidelity counterfactual healthy images can further enhance the interpretability of the segmentation models, as well as provide a reliable baseline for evaluating the extent of anomalies and supporting clinical decision-making. Notably, the unsupervised anomaly detection model is trained purely on healthy images, eliminating the need for anomalous training samples and pixel-level annotations. We validate the proposed approach on carotid US, brain MRI, and liver CT datasets. The experimental results demonstrate that the proposed framework outperforms existing state-of-the-art unsupervised anomaly detection methods, achieving performance comparable to fully supervised segmentation models in the US dataset. Additionally, ablation studies underline the importance of hyperparameter selection for Synomaly noise and the effectiveness of the multi-stage diffusion process in enhancing model performance. Code: https://github.com/yuan-12138/Synomaly.
keywords:
Ultrasound Image Analysis, Unsupervised Anomaly Detection, Diffusion Model, Medical Image Analysis, Carotid imaging
1 Introduction
Ultrasound (US) plays a crucial role in daily medical practice. It is commonly used for the screening and early assessment of internal organs and vasculature. Compared to computed tomography (CT) and magnetic resonance imaging (MRI), US is more frequently utilized for routine examinations. CT, although efficient in providing high-quality images, involves radiation exposure, which makes it less suitable for regular check-ups. MRI, while providing detailed soft tissue contrast, is often cost- and time-intensive for routine use. Consequently, physicians typically reserve CT and MRI for cases where a patient has already exhibited specific symptoms or conditions. As a result, CT and MRI are more commonly associated with the imaging of diseased cases. In contrast, US is radiation-free, cost-effective, and portable, making it an ideal modality for regular monitoring. Therefore, the US images collected are often from non-diseased individuals, making it as an ideal modality to learn the physiology of healthy tissues.
However, the inherent challenges of low reproducibility and poor image quality in US imaging make it difficult to gather large amounts of diseased data for training of deep neural networks. This challenge is particularly evident in conditions like stroke, a leading cause of mortality worldwide, with ischaemic stroke accounting for about 65.3% of all cases [16]. This type of stroke is caused by the blockage of a blood vessel supplying the brain, and carotid atherosclerosis is a common underlying cause [22]. US is frequently used for the detection and monitoring of carotid plaques. One important indicator to assess the risk of stroke for atherosclerosis patients is total plaque volume [1]. One pre-requisite for such measurement is an accurate segmentation of the carotid plaque in US images. However, unlike healthy anatomy, plaques vary significantly in shape and location, making their detection more challenging. This variability means that high-quality datasets are crucial for training segmentation models to reliably identify plaques. Without such datasets, models may struggle to generalize and correctly segment plaques in diverse cases. Because of ethical and privacy issues, it is not an easy practice to gather large amount of diseased US data. To tackle such challenge, unsupervised anomaly detection strategy appears to be an effective solution.
Similar to how humans learn about diseased tissues, unsupervised anomaly detection method tends to distinguish the anomaly tissue from the normal anatomy by learning from healthy structures [17]. Generative models are commonly utilized to generate synthetic counterfactual healthy samples based on diseased images. Since the model is purely trained on healthy data, the need for pixel-level labels is eliminated. Considering the high noise ratio and low contrast of US compared to CT and MRI, the labelling process for US sometimes introduces unavoidable ambiguities, thus complicating the fully-supervised training process [15]. However, such drawback does not concern the unsupervised anomaly detection method. Furthermore, it also provides the possibility to detect unseen or rare anomaly cases, which are not well represented in the dataset for fully-supervised models. Nonetheless, unsupervised anomaly detection is less common in US due to the presence of speckle noise, shadows, and acoustic artifacts, which make it challenging to differentiate anomalies from healthy tissues. Such complexity of US images necessitates the development of more advanced generative models.
Diffusion models have proven effective in modeling complex data distributions, resulting in higher image sample quality compared to generative adversarial networks (GANs) and variational autoencoders (VAEs) [12, 10]. These advantages are particularly valuable for unsupervised anomaly segmentation, especially in cases of small datasets, leading to growing interest in applying diffusion models to this application domain [27]. Most recent works have also reported initial attempts of applying diffusion models for unsupervised anomaly detection in US [34, 32], specifically exploring their effectiveness for detecting anomalies in fetal brain US. The performance of diffusion model is closely linked to the choice of corrupting noise, especially for medical image anomaly detection [26]. Superior performance over naive pixel-wise Gaussian noise has been reported when applying Simplex noise [42] and Coarse noise [25] for anomaly detection of brain MRI.
In order to develop a robust unsupervised anomaly detection model for US images, in this work, we propose Synomaly noise, which stands for synthetic anomaly noise based on statistical disease morphology prior. Different from the aforementioned noise functions, Synomaly noise is specifically designed for unsupervised medical anomaly detection by corrupting healthy images with synthetic anomalies during training. This design of noise can result in a better erasure of anomalies in the images. To further suppress the loss of detailed information during noising process, a multi-stage diffusion process is proposed for inference. The detailed contributions are summarized as follows:
1.
We introduce the Synomaly noise function to enhance the effectiveness of anomaly detection. This noise is applied to healthy training images to simulate the appearance of real anomaly tissues after the noising process. By exposing the model to these synthetic anomalies during training, the effectiveness of generating healthy images from diseased images is significantly improved, tailored to individual diseases.
2.
We propose a multi-stage diffusion process in the inference phase designed to preserve healthy tissue details while precisely isolating and removing anomalous disease tissue. Instead of executing the diffusion process with a large number of noising and denoising steps in one single stage, the generation is divided into successive diffusion stages, each with fewer noising and denoising steps. Through such gradual and controlled transformation, the model can preserve finer details in the reconstructed images, thus increasing the accuracy in anomaly detection tasks.
3.
This is the first work to apply a diffusion model for unsupervised anomaly detection in carotid US images. We demonstrate the effectiveness of the proposed Synomaly noise and multi-stage diffusion process on a US carotid plaque dataset. To further validate the performance of the proposed approach, we conducted experiments on public brain MRI and liver CT datasets, achieving superior performance over state-of-the-art methods in detecting brain tumors and liver diseases, respectively.
The results of the three most common medical image modalities (US, CT, and MRI) demonstrate the potential of the proposed method in wide application. To facilitate the reproducibility of this work, the code will be released on this website upon acceptance.
2 Related Work
Anomaly segmentation is based on the principle that a network trained solely on normal images will struggle to reconstruct regions containing abnormalities. The model learns to accurately recreate typical patterns of healthy tissue but fails when facing with abnormal areas. During inference, the model attempts to reconstruct the entire input image, but the anomalous regions remain poorly reconstructed due to the model’s lack of exposure to such data. By subtracting the reconstructed image from the original input image, the differences—representing the anomalies—can be highlighted, which enables the pixel-wise identification of abnormal regions [4]. From a methodological perspective, anomaly segmentation approaches can also be grouped into two main categories: generative autoencoders and denoising models. The following sections will provide an overview of related works in these two categories.
2.1 Unsupervised Anomaly Detection Using Autoencoders
Autoencoders (AEs), since their introduction, have been widely used for unsupervised representation learning. The principles behind AE-based unsupervised anomaly detection/segmentation is to learn a latent space distribution that represents healthy anatomies, identifying abnormalities as outliers. Consequently, the reconstructed images preserve the healthy regions while altering the anomalous parts, making it easier to detect and segment the abnormalities. Different variants of AEs have been applied to different anomaly detection scenarios. [40] implemented a conditional VAE to perform anomaly segmentation in brain MRI, while [35] utilized a Bayesian convolutional AE to locate the abnormalities in CT. By combining VAE with a self-supervised image representation learning network, [29] proposed SSL-AnoVAE to tackle the anomaly staging and segmentation of retinal images. Deviated from the reconstruction training approaches, [44] adopted an inpainting strategy to train a context-encoding VAE for anomaly detection in brain MRI. In order to increase the ability of AEs in modeling complex high-dimensional data distributions, adversarial strategies have been introduced [21]. [5] integrated GAN with VAE and proposed AnoVAEGAN to perform anomaly detection for brain MRI.
Apart from reconstruction-based approaches, restoration-based methods offer an alternative. Instead of directly passing input images through AE network to generate counterfactual reconstructions, restoration methods focus on finding the closest representation of the abnormal image in the learned latent space of normal images. These methods project the abnormal image onto the normal manifold and transform it into its closest healthy version. Anomalies are then detected by comparing the original input with the restored image. [39] proposed AnoGAN, where a GAN is trained on healthy images to model the latent manifold for normal images. During inference, a random latent point is sampled and iteratively moved towards the closest counterpart on the normal manifold using a gradient decent method. In an extended work, [38] implemented an encoder network to replace the time-consuming iterative restoration process. [11] trained a VAE to model the normative data distribution. The restoration process of anomaly samples is performed through an iterative gradient ascent method by Maximum-A-Posteriori estimation.
2.2 Unsupervised Anomaly Detection Using Denoising Models
Recent advances in diffusion models have demonstrated superior generative performance across various medical imaging modalities [13, 43, 14, 24]. This enhanced performance has also proven beneficial for unsupervised anomaly detection in medical imaging, enabling more accurate identification of abnormalities [27, 28]. [42] proposed AnoDDPM, where a Denoising Diffusion Probabilistic Model (DDPM) is utilized to reverse the diffusion process and generate counterfactual images. A partial diffusion process is introduced to better conserve the fine details of the input images and further increase the sample time efficiency during inference. They also replaced the common Gaussian noise with Simplex noise for image corruption and achieved better results when testing on brain MRI images. To avoid the accumulated errors during iterative denoising process and increase the sample efficiency of the diffusion process, [36] implemented latent diffusion model for anomaly detection in brain MRI and CT. The noising and denoising process is then executed in the latent space rather than directly on the images, which can significantly lower the dimensionality and improve efficiency. [41] utilized a Denoising Diffusion Implicit Model for medical anomaly detection, which enables a much faster sampling than DDPMs. The diffusion model is trained on both healthy and diseased images, and during inference, the generation of healthy counterparts for diseased images is guided by a pretrained classifier. Instead of providing direct guidance during the denoising process, [8] introduced THOR to provide implicit guidances through intermediate anomaly maps. This approach allows the model to focus on potential anomaly regions while preserving more details in the healthy areas, thereby reducing the false positive rate in anomaly detection. Similarly, [19] combined DDPM with ACAT [18], a saliency map that can approximately capture the pathological regions in the input image. At each denoising step, the regions identified by the ACAT saliency map are processed separately enabling targeted modification of potential diseased areas while preserving healthy regions. As a result, they demonstrated enhanced erasure performance of the anomalous regions in experiments on brain MRI and CT datasets. Considering the high computational burden of the diffusion process for high-resolution images like MRI, [6] proposed a patch-based diffusion model, which reconstructs input images patchwise, thereby reducing network complexity.
Given the nature of diffusion process, the corruption of input images will inevitably lead to the loss of details of the original images. However, too little corruption of the input images may also lead to unsatisfactory anomaly erasure performance. To tackle such trade-off, [25] proposed denoising AE (DAE) that applies a single instance of Coarse noise to the input image, rather than sequentially adding noise. This one-time noise corruption alters the original image to resemble an anomalous version. By training the DAE to recover the corrupted images, the model can generate counterfactual normal images from pathological ones during inference. In a follow-up work, [26] highlighted the importance of the noise function in diffusion models and DAEs for anomaly segmentation, showing that the choice of noise significantly impacts the models’ ability to distinguish between normal and anomalous regions. Inspired by this insight, [30] employed a masked noising strategy, where only a selected region of the input image is corrupted. During inference, the DAE is iteratively applied until the generated anomaly mask converges.
3 Method
The performance of diffusion models in unsupervised anomaly detection is closely related to the design of the noising function. An effective noise function should efficiently corrupt the original image such that, after the noising process, the anomalous regions become indistinguishable from the normal ones. This allows the model to replace anomalies with normal tissues during the denoising process. However, the classic noise functions face challenges with larger or particularly distinct anomalies, such as very bright or dark regions. In such cases, many noise steps may be required to sufficiently blur the anomaly and make it indistinguishable from the surrounding healthy tissue. This extensive noise addition can lead to significant loss of structural image information, resulting in blurred and inaccurate image reconstruction. This introduces a precision-recall trade-off: Using fewer noise steps preserves structural integrity and achieves high reconstruction accuracy for healthy tissues, but it increases the risk of failing to detect distinct anomalies, leading to a higher rate of false negatives - characterized by high precision and low recall. On the other hand, using more noise steps enhances the likelihood of detecting anomalies, but it also raises the possibility of false positives, where healthy tissues are incorrectly identified as anomalies - characterized by low precision and high recall.
To tackle these challenges, we propose Synomaly noise. The core idea is to corrupt healthy images by introducing synthetic anomalous regions during training, simulating the appearance of noised anomalous regions, as shown in Fig. 1(a). This allows the model to refine its denoising strategy specifically for reconstructing normal tissue from noised regions that resemble pathological anomalies. During inference, standard Gaussian noise is applied to the images, causing the abnormal regions to resemble the noised artificial anomalies encountered during training. This similarity enables the model to efficiently generate counterfactual healthy tissues and effectively replace anomalies with normal tissue during the denoising process.
To further improve the conservation of fine details during diffusion process, we employ an iterative anomaly erasure approach, as shown in Fig. 1(b). In this method, images undergo multiple rounds of the diffusion process, with anomalies being progressively removed in each iteration until the anomaly map converges. This iterative process is especially beneficial in cases where anomalies are large, as larger noising steps are typically required to fully obscure the abnormal regions. By splitting these larger noising steps into smaller, incremental steps, the model reduces the risk of losing fine details that often occurs when excessive noise is applied at once.
3.1 Diffusion Models
The diffusion model consists of two parts: a forward (noising) process and a reverse (denoising) process. In the forward process, noise is progressively added to a data sample over a sequence of noise steps, gradually corrupting the data and transforming it into a noise-like distribution. The corrupted image can be described as:
where represents the original image, is the corrupted image at step , depicts the noise function, while acts as a schedule, controlling the strength of the corruption at each step . Its value starts from and ends with . represents the total noising steps, which we set as in our setup. This schedule can be a linear schedule as proposed by [23] or a cosine schedule as introduced by [33].
The reverse diffusion process aims to invert the forward process by gradually removing the added noise from the data, reconstructing the original data from the noisy state. It can be achieved by using a denoising network with trainable parameter , normally a U-shape architecture [37]. [23] found that instead of directly predicting the denoised image at each step, predicting the noise term () works the best among several possible objectives. Hence, the optimization function is defined as:
After the training, given a corrupted image or a complete noise we can generate a novel image by iteratively performing the denoising sampling based on DDPM [23]:
with
3.2 Synomaly Noise
Synomaly noise stands for synthetic anomaly noise. Instead of merely adding noise until anomalies blend into healthy tissues, Synomaly strategically integrates synthetic anomalies into the noise model during training. The core concept is to train the diffusion model to remove synthetic anomalies in healthy images that mimic real-world scenarios, thereby improving the model’s ability to detect and remove actual anomalies during inference. The key advantage of Synomaly in comparison to the other noise functions is its ability to utilize disease-specific information (morphology or appearance of the disease) during model training, which greatly enhances anomaly detection performance based on our experiments.
Fig. 2 illustrates the process of Synomaly noise generation. In the first step, a pure Gaussian noise is created, matching the dimensions of the input image. This Gaussian noise will also serve as the Gaussian background noise of Synomaly noise. Next, the shapes of the synthetic anomalies are generated by applying Gaussian blurring to the initial noise, followed by thresholding. Here, the variance of Gaussian blurring () as well as the threshold value () are predefined as model training parameters and control the size and number of synthetic anomalies. Specifically, increasing the blurring variance () results in larger fragments of anomalies, while higher threshold values () reduce the number and size of anomalies by excluding lower Gaussian values. To further refine the simulation of real anomalies, an optional mask can be applied to constrain the anomaly shapes to relevant regions.
Once the synthetic anomaly shapes are defined, synthetic anomalies are added to the Gaussian background noise at their respective locations. A predefined anomaly direction determines whether the anomalies typically appear brighter or darker than the surrounding healthy tissue and a predefined intensity offset () dictates the distinctiveness of the synthetic anomalies. Additionally, a random value () between 0 and 1 is added to the intensity of these synthetic anomalies to enhance robustness against variations in actual anomaly brightness.
In summary, the tunable parameters include blurring variance , threshold value , anomaly direction and intensity offset . Guided by these parameters to reflect realistic medical conditions, the resulting synthetic anomalies vary randomly in number, shape, size, intensity and location, as they are all derived from pure Gaussian noise.
3.3 Multi-stage Diffusion Process at Inference
Most of the existing works on diffusion models for anomaly detection typically suggest performing the diffusion process only once during inference, which is termed as “single-stage” approach. However, this approach is usually insufficient when dealing with anomalies that are unique in shape or intensity. Often, a single-stage cycle cannot fully remove anomalies, especially when many noise steps are required to achieve good anomaly detection. While increasing the number of noise steps can improve anomaly removal, it can also degrade structural information and eventually lead to blurred and inaccurately reconstructed images.
Therefore, we propose a “multi-stage diffusion process”. Instead of applying many noise steps in a single inference stage, the multi-stage approach repeatedly applies a few noise steps followed by denoising across multiple inference stages. As depicted in Fig. 3(a), starting with the original input image, the process begins by adding noise for a predefined number of noise steps, followed by denoising. Then, the denoised image is fused with the original image to have the image with reduced abnormality for the next stage. This fusion process is detailed in Fig. 3(b), where the computed anomaly mask in the current stage is used to provide guidance. This process is repeated over multiple stages until the generated anomaly map converges. To prevent excessive iterations, the maximum number of iterations is set to five.
This approach offers several benefits. By using fewer noise steps at each stage, the structural integrity of the image is better preserved, leading to more accurate reconstructions. Additionally, performing multiple inference stages with sequential denoising allows even distinct anomalies to be progressively removed and detected more effectively. Since different noise steps model different image features [23] and the goal of medical anomaly detection is to focus on finer details without excessively corrupting the entire image, fewer noise steps are particularly relevant for detecting subtle anomalous regions.
To further reduce the false positive rate in the generated anomaly mask, we propose a “masked fusion” strategy, as shown in Fig. 3(b). At each diffusion stage, rather than directly passing the generated image to the next cycle of noising and denoising, we mask the image by the anomaly mask produced at that stage and fuse it with the original input image. This approach preserves the “normal” regions from the input image, while maintaining the reduced anomalies at each stage. As a result, the quality of the generated healthy image is enhanced, effectively reducing the false positive detection rate. The detailed process is summarized in Algorithm 1.
4 Experiments
In this section, we evaluate the performance of the proposed model using various datasets. Since the US carotid plaque dataset is collected in-house, we additionally present evaluation results on publicly available MRI and CT datasets to demonstrate the superiority and effectiveness of the proposed method.
4.1 Datasets
In this work, three datasets are used, namely, a brain MRI dataset, a liver CT dataset, and a carotid US dataset. The details of each dataset are provided below.
4.1.1 Brain MRI: BraTS Dataset
The BraTS2023 dataset originates from the international Brain Tumor Segmentation (BraTS) Challenge [2, 3, 31], which is a well-established benchmark in the field of medical imaging. It comprises 3D brain MRI scans from patients diagnosed with brain tumors, providing a rich and diverse set of data for developing and evaluating segmentation models. The pixel-wise annotations classify tissues into the four categories healthy tissue, GD-enhancing tumor (ET), peritumoral edematous/invaded tissue (ED) and necrotic tumor core (NCR). For the purposes of this work, all three tumor-related classes (ET, ED, and NCR) are grouped into a single “anomalous” category, forming two classes: healthy and anomalous tissue.
Since the focus is on tumor detection in 2D images, only axial brain slices are considered in this work and the 2D tumor ground truth segmentation is extracted from the annotated 3D classification volumes. Given that ET, ED and NCR are most commonly found in the cerebral hemispheres, the brain MRI dataset is further refined to include only the central slices. For anomaly detection, the FLAIR sequence is used. All selected slices are resized to pixels to have unified dimensions. anomalous slices are randomly selected as the test set. In total, the dataset consists of healthy training samples and anomalous test samples. To help intuitive understanding, a few representative samples are visualized in Fig. 4(b).
4.1.2 Liver CT: LiTS Dataset
The Liver Tumor Segmentation Benchmark (LiTS) dataset [9] includes abdominal CT volumes, each with detailed annotations of the liver and its tumors. The number of tumors per volume ranges from to , containing both primary and secondary tumors of diverse sizes and appearances. For the purposes of this work, all tumors are categorized as “anomalous” class.
Each 3D abdominal CT volume contains varying axial slices, ranging from and . To prepare the data for model training and evaluation, the annotated segmentations are used to classify slices into healthy and anomalous cases. Further refinement is performed using size thresholding to isolate the relevant slices and all images are resized to pixels. In total, the CT dataset has healthy training images and anomalous test images. A set of representative samples is shown in Fig. 4(c). These examples highlight the complexity of the anomaly detection task, as the abdominal region exhibits various appearances and the liver itself can differ in size, shape and location. The relatively small ratio between the anomaly and the chest volume can further cause class imbalance problems. These factors make it particularly challenging to accurately segment liver tumors, while preserving healthy tissue. This high diversity and complexity of the LiTS dataset make it a great benchmark for assessing the performance and generalizability of the proposed anomaly detection approaches.
4.1.3 Carotid US Dataset
The carotid artery US dataset includes US scans from healthy patients and patients with carotid plaques. The US images in the dataset are all captured from the cross-sectional view of the carotid artery. The acquisition was performed within the Institutional Review Board Approval by the Ethical Commission of the Technical University of Munich (reference number 244/19 S). Due to the nature of US imaging, B-mode appearance is affected significantly by contact point, force, angle, and US machine setting. So, to alleviate the effects caused by the changing image appearance, we crop the US image around the target anatomy, i.e., the carotid artery, to ensure the network only focuses on disease-cased image variations. In this study, we applied existing vessel tracking and segmentation methods to generate vessel masks, which are then used to crop the original full-view US images. Fig. 4(a) presents the pipeline for vessel tracking and cropping procedure. To account for potential imperfections in vessel segmentation and to enhance model robustness, a small degree of randomness regarding the position and size of the vessel within the image is introduced in the cropping process, which also creates some variability in the training data.
The ground truth annotations for plaques in the anomalous carotid artery samples are manually annotated under the supervision of our clinical experts specialized in vascular disease. All such annotations of vessels and anomalies were performed using the ImFusion Suite111https://www.imfusion.com/products/imfusion-suite. All US images are resized to pixels to standardize input dimensions. In total, the final dataset contains healthy training images and anomalous test images. The examples can be found in Fig. 4(d).
| Method | Noise steps () | Kernel size () | Threshold () | ||||||
| MRI | CT | US | MRI | CT | US | MRI | CT | US | |
| VAE | 13 | 13 | 13 | 0.2 | 0.35 | 0.3 | |||
| DDPM+Gaussian | 300 | 300 | 800 | 15 | 15 | 15 | 0.15 | 0.2 | 0.3 |
| DDPM+Corase | 700 | 800 | 800 | 1 | 5 | 21 | 0.15 | 0.1 | 0.1 |
| DDPM+Pyramid | 350 | 250 | 800 | 13 | 13 | 5 | 0.2 | 0.2 | 0.35 |
| DDPM+Simplex | 200 | 200 | 700 | 1 | 1 | 5 | 0.2 | 0.25 | 0.45 |
| AutoDDPM | 200 | 400 | 700 | 9 | 3 | 13 | 0.1 | 0.2 | 0.1 |
| DAE | 11 | 9 | 3 | 0.1 | 0.25 | 0.1 | |||
| DDPM+Synomaly+Single-stage | 200 | 200 | 400 | 15 | 11 | 11 | 0.15 | 0.25 | 0.25 |
| DDPM+Synomaly+Multi-stage | 15 | 5 | 15 | 0.2 | 0.2 | 0.3 | |||
Noise steps at each stage.
4.2 Inference Hyperparameters
The number of inference noise steps plays a significant role in balancing accurate image reconstruction and effective anomaly detection. As discussed earlier, methods like optimizing the noise function can be employed to address this trade-off, but it remains essential to find the optimal balance between anomaly removal and image preservation for specific tasks. Therefore, it is a critical hyperparameter dominating the anomaly detection results.
Directly computing the pixel-wise difference between the original and generated images is not always the most suitable approach due to potential imperfections in pixel-wise image reconstruction. Moreover, in the context of preventive medical anomaly detection, the focus lays on identifying regions of potential anomalies rather than individual pixels, so that medical professionals can further evaluate these areas. Therefore, Gaussian blurring () is applied to both the input and generated images before calculating their differences. The size of the Gaussian kernel is denoted as , denoting a second hyperparameter that influences the anomaly segmentation performance.
In addition to the inference noise steps and Gaussian kernel size , a third crucial parameter during inference is the anomaly threshold . After the difference map is created by calculating the pixel-wise residuals between the blurred original image and blurred generated image , anomalies are identified by applying this anomaly threshold to the difference map, flagging regions with differences above this threshold as anomalies.
In summary, the inference noise steps for partial diffusion (), the Gaussian kernel size for blurring (), and the anomaly threshold for anomaly selection () are the three hyperparameters that need to be carefully tuned for each dataset and diffusion noise function. To ensure a fair comparison between different models, the best inference hyperparameter is found for each model, representing the upper limit of performance.
4.3 Training Details
The performance of the proposed anomaly detection method is compared with six state-of-the-art (SOTA) methods: classic VAE, DDPM using Gaussian noise function [23], DDPM using Coarse noise function [26], DDPM using Pyramid noise function [20], DDPM using Simplex noise function [42], AutoDDPM [7], and DAE [25]. Fully supervised UNets [37] were also trained for each dataset as a reference. Half of the anomalous data of each dataset is selected as UNet training data.
The VAE model and DAE model are trained for epochs using Adam optimizer with an exponentially decreasing learning rate from to by . All the DDPM models are implemented based on the popular guided diffusion approach by OpenAI [12]. For the Simplex noise function, the noising parameters were set with an octave of 6, a persistence of 0.8 and a frequency of 64, as recommended by the original authors. For Coarse noise, the noising parameters were configured with a noise resolution of 16 and a noise standard deviation of 0.2, also aligning with the established literature. All the DDPMs are trained using Adam optimizer with a fixed learning rate of . The models were trained for 3000 epochs on the brain MRI and liver CT datasets, and for 2000 epochs on the carotid US dataset. All the models were trained using Nvidia A40 GPU.
4.4 Results on Brain MRI
For the anomaly detection in brain MRI images, the Synomaly noise function parameters are configured with a blurring variance of 3, a threshold value of 175, an anomaly direction of 1 and an intensity offset of 0.5. These Synomaly parameters were determined through a combination of known disease morphology and experimental validation. To ensure a fair comparison between methods, the optimal inference hyperparameters for each method were determined through a comprehensive grid search. Tab. 1 presents the best hyperparameter combinations for each method. Notably, for VAE and DAE, since no diffusion process is involved, the “noise steps” field is left blank. “DDPM+Synomaly+Multi-stage” depicts the case where the proposed Synomaly noise is used to train the DDPM and the multi-stage diffusion process is utilized during inference. The “noise steps” field for “DDPM+Synomaly+Multi-stage” indicates the noise steps at each stage of the iterative process.
Using the best parameters identified, each model was tested on 1000 anomalous images. The evaluation metrics include the Dice, Precision, and Recall score, to ensure an objective and comprehensive comparison of different approaches. The anomaly segmentation results are shown in Tab. 2. It is evident that the proposed Synomaly noise function consistently outperforms all other unsupervised methods across all metrics. Fig. 5 shows the visual results of the reconstructed images and anomaly segmentation maps using different unsupervised methods. DDPMs trained using Coarse noise and Simplex noise functions partially erase the anomalous regions, but distinct borders of the brain tumors remain in the reconstructed images. For AutoDDPM and DAE, the bright anomalous area is slightly dimmed but not completely removed. However, changes in these areas can still be identified with a proper threshold. The VAE reconstruction is anomaly-free but appears blurry, losing important brain details. Meanwhile, DDPMs trained with Gaussian noise and Pyramid noise fully remove the round anomaly in the top area of the image but fail to accurately reconstruct the lateral ventricle. In contrast, the image reconstructed using the proposed method shows a more realistic reconstruction, particularly for the lateral ventricle, demonstrating its superior performance in terms of generative fidelity. The proposed approach delivers the best anomaly segmentation mask compared to the other SOTAs.
4.5 Results on Liver CT
The configuration of Synomaly noise function for liver CT anomaly detection task is set as a blurring variance of 3, a threshold value of 175, an anomaly direction of -1 and an intensity offset of 1. Note that for this specific use case, the anomaly detection direction was set to -1 instead of 1, since the liver tumors appear darker than the surrounding healthy tissue, requiring an inversion of the detection parameter.
The best inference hyperparameters are determined for each method by applying grid search. The anomaly segmentation performance is summarized in Tab. 2. In this scenario, the VAE completely fails to generate plausible counterfactual samples, as the complexity of CT images exceeds the generative capacity of the VAE model. As shown in Fig. 5, the VAE-generated images are of low resolution with jagged borders. The DDPMs trained with Gaussian and Pyramid noise function can effectively erase the abnormalities but they also corrupt other parts of the image, resulting in a high false-positive detection rate. The AutoDDPM and DAE cannot fully remove the anomalies, but with a proper threshold, the anomalous regions can still be highlighted. The DDPM trained with the Simplex noise function demonstrated relatively good generation results, with liver tumors almost fully removed. The best performance was observed with Synomaly noise plus the multi-stage diffusion process, which achieved the highest scores across all metrics. The visual results also confirm its superior performance, as the anomalies are completely erased without any trace. When comparing the performance of unsupervised anomaly segmentation methods to fully supervised segmentation models, a significant gap remains in this scenario. This can be attributed to the complexity of abdominal CT images, where the generative model must reconstruct not only the liver but also all other organs in the abdomen, making accurate segmentation considerably more challenging.
4.6 Results on Carotid US
The Synomaly noise function in this scenario was defined with parameters based on known disease morphology and experimental validation, which resulted in a blurring variance of 7, a threshold value of 150, an anomaly direction of 1 and an intensity offset of 0.5. To identify the best inference parameters for each model, a comprehensive grid search was conducted across the parameters noise steps (), Gaussian kernel size () and anomaly threshold (). Tab. 1 presents the optimal parameters for each method.
The anomaly segmentation results are summarized in Tab. 2. It is evident that Synomaly noise significantly outperforms all other unsupervised methods. Interestingly, the best performance apart from the proposed approach in this scenario is achieved by the VAE, with an average Dice score of . This can be attributed to the fact that, compared to brain MRI and liver CT, the structure of the carotid vessel is relatively simple, allowing the VAE to reconstruct a plausible plaque-free artery from a diseased one. Additionally, carotid plaques are more distinct and larger in size, which makes removing such anomalies with traditional SOTA noise functions challenging. Large noising steps are required, resulting in severe corruption of the image and significant information loss. Due to the similarity in textures between plaques and surrounding tissues, noising further blurs the distinction between these structures. As observed in the fifth row of Fig. 5, generated vessels often incorrectly interpret the inner border of the original diseased vessel as the outer border, leading to reconstructions of smaller-than-actual arteries. The DAE fails to effectively remove the plaques partly because the added noise during training does not sufficiently represent the anomalies. Consequently, during inference, the network encounters situations that it was not exposed to during training. In contrast, Synomaly noise effectively identifies and removes the plaques within the artery. The image generated using the proposed approach demonstrates the best performance among all methods evaluated. As illustrated in the last subplot of the fifth row of Fig. 5, the plaques are completely removed, and the artery’s borders are accurately reconstructed.
Compared to the fully supervised method, the proposed unsupervised approach is only points lower in Dice score and even achieves a higher Recall. Considering the small size of the carotid plaque dataset compared to brain MRI and liver CT, the reduced segmentation performance of fully supervised method can be attributed to the scarcity of diseased data. However, this limitation reflects the typical reality of US imaging, where large diseased datasets are difficult to obtain, while healthy datasets are easily accessible. Notably, the healthy US dataset was collected using a Siemens machine using default “Carotid” acquisition parameter setting, whereas the anomalous dataset was recorded using a Philips machine with a sonographer-defined parameter setting. Despite the existing domain gap, the proposed anomaly detection model still demonstrates comparable performance to the fully supervised method, highlighting its robustness and effectiveness in practice.
4.7 Ablation Study
In this section, we conduct ablation studies on the hyperparameters of Synomaly noise and the multi-stage diffusion process. These experiments aim to provide readers with a clearer understanding of how these hyperparameters influence the final anomaly segmentation performance and highlight the importance of employing the multi-stage diffusion process.
4.7.1 Hyperparameter Ablation for Synomaly Noise
In this ablation study, we first examine the influence of the hyperparameters of Synomaly noise, as described in Sec. 3.2. The performance of Synomaly noise is closely linked to its ability to represent the target anomalies. If the Synomaly noise generated during training resembles the real abnormalities encountered during inference, the trained model can more effectively erase the diseased areas during testing. The morphology of the generated Synomaly noise is controlled by two parameters: the blurring variance () and the threshold value ().
To evaluate the influence of these parameters, several models were trained with different values of and on the carotid US dataset. The generated noise is normalized to the range [0, 255], and consequently, the threshold values are also within this range. First, we kept fixed at 150 and varied from 1 to 11. The results are presented in Fig. 6(a), which shows that the best performance was achieved with a blurring variance of 7. Similarly, for , we conducted experiments by fixing at 7 while varying from 90 to 230. The results, as shown in Fig. 6(b), indicate that the best performance was achieved when setting to 150. Some examples of the generated Synomaly noise are displayed below the plots. The value of controls the size of the synthetic anomaly regions, while determines the number of generated artificial anomaly areas. A smaller value results in smaller synthetic anomaly regions, whereas a lower value leads to a higher number of generated anomaly areas. However, in some cases, when choosing a smaller value, these generated regions sometimes overlap with each other, leading to fewer but larger anomalous areas. For the case of US plaque detection, the optimal choice is , , since this combination can best resemble the carotid plaque morphology based on our experiments.
4.7.2 Ablation for Multi-stage Diffusion Process
To validate the effectiveness of the proposed multi-stage diffusion process, we conduct an ablation experiment to compare the anomaly segmentation performance with and without multi-stage diffusion. The DDPM was trained using Synomaly noise, and during inference, we tested the model’s performance with a single-stage diffusion approach. To ensure a fair comparison, the optimal inference hyperparameters for single-stage diffusion with Synomaly noise were determined using a grid search and are reported in Tab. 1, denoted as “Synomaly+Single-stage”. In general, the number of noise steps at each stage for multi-stage inference is much smaller than the number of noise steps for single stage, which helps to preserve finer details. The quantitative results can be found in Tab. 3 under the “Synomaly+Single-stage”, and the visual results can be found in Fig. 7 on the right part.
| Metric | Brain MRI | Liver CT | Carotid US |
| Synomaly+Single-stage | |||
| DSC | |||
| Recall | |||
| Synomaly+Multi-stage w/o Masked Fusion | |||
| DSC | |||
| Recall | |||
| Synomaly+Multi-stage w/ Masked Fusion | |||
| DSC | |||
| Recall | |||
For the case of brain MRI dataset, the single-stage DDPM trained using Synomaly noise successfully erases the round anomaly but struggles with generating a realistic lateral ventricle, as shown in Fig. 7(e). The results using the multi-stage diffusion process show a more realistic reconstruction, particularly for the lateral ventricle, compared to the single-stage process. The detailed generation procedure of using multi-stage diffusion process can be found in Fig. 3(a). The anomalies are gradually removed, especially the abnormalities that appear in the area of lateral ventricle. As depicted in Tab. 3, the multi-stage diffusion process did not significantly improve anomaly segmentation in this case, showing similar performance compared to the single-stage approach. This is partly because the anomalies in FLAIR-weighted MRI are relatively clear, allowing the model to achieve satisfactory results with just a single iteration, making multiple diffusion iterations unnecessary. However, the comparison between Fig. 7(a) and Fig. 7(e) demonstrates that images generated using the multi-stage diffusion process are more realistic compared to those produced with a single diffusion step.
For the case of liver CT, the multi-stage diffusion process has successfully increased the Dice by around 0.03 compared to the single-stage approach. The visual results also support this observation. As depicted in Fig. 7(f), the single-stage Synomaly method mostly removes the anomalies, but faint traces remain visible upon close inspection. In contrast, the multi-stage approach completely eliminates all the anomalies without a trace. Fig. 6(b) shows how the abnormalities are progressively removed in each iteration of multi-stage diffusion process.
For US carotid plaque anomaly segmentation, implementing the multi-stage diffusion process has improved the Dice score by approximately compared to the single-stage approach. The two visual examples in Fig. 7(g) and Fig. 7(h) also demonstrate that a single diffusion process cannot fully erase the anomalous regions. While the plaque in the upper part of the vessel is successfully removed, the larger plaque at the bottom of the vessel remains in the generated image. After applying multi-stage diffusion process, all the plaques are fully removed. Fig. 7(c) shows the generated images at each stage, clearly illustrating how the plaques are iteratively removed from the original image. Generally, distinct and large anomalies require more iterations to be completely erased. Fig. 7(d) showcases another example involving five iterations, where the distinct plaque at the bottom of the vessel lumen is progressively removed over five iterations, while the smaller plaque at the top is mostly removed in four iterations.
Additionally, we evaluated the performance without using the masked fusion technique, where the generated image from the previous iteration was directly used as input for the next iteration of the diffusion model. The results, labeled as “Synomaly+Multi-stage w/o Masked Fusion”, are shown in Tab. 3. Compared to the results of applying masked fusion technique, labeled as “Synomaly+Multi-stage w/ Masked Fusion”, the masked fusion led to an increase of approximately 0.02 in Dice score across all datasets. Without masked fusion, the non-anomalous parts of the image also undergo multiple diffusion processes, which may introduce changes to healthy regions. By using the masked fusion technique, the original texture in non-diseased areas is preserved, thereby enhancing anomaly segmentation accuracy.
5 Conclusion and Discussion
In this paper, we presented a novel framework for unsupervised anomaly detection in US images using Synomaly noise and a multi-stage diffusion process. Synomaly noise was designed to add synthetic anomalies to healthy US images during training, enabling the model to effectively remove anomalies during inference. The multi-stage diffusion process was introduced to incrementally denoise the images, thereby preserving fine details and generating more realistic reconstructions.
The experiments on carotid US, brain MRI, and liver CT datasets demonstrated the efficacy of the proposed approach. For the carotid US dataset, the proposed method achieved performance comparable to fully supervised segmentation models, demonstrating its strong anomaly removal capabilities even without large amounts of labeled data. The proposed method improved unsupervised anomaly segmentation results by 0.23 in the carotid US dataset. The brain MRI experiments showed an increase of approximately 0.1 in Dice score with the proposed approach, while the liver CT dataset demonstrated an increase of 0.04, showing improved unsupervised anomaly segmentation accuracy. Additionally, the ablation studies highlighted the importance of configuring reasonable hyperparameters for Synomaly noise, as its performance is closely tied to its ability to resemble the target anomalies. We also demonstrated the effectiveness of the multi-stage diffusion process and the masked fusion technique in improving segmentation quality, both of which contributed to enhanced model performance.
Overall, the proposed approach outperformed existing state-of-the-art methods for unsupervised anomaly detection, significantly narrowing the gap between unsupervised and fully supervised approaches. Despite challenges such as variability in acquisition settings and machine types, the proposed method exhibited robustness, generalizability, and promising potential for clinical applications. Moreover, the proposed framework provides an opportunity for doctors to incorporate their prior knowledge of the disease’s appearance, thus enhancing unsupervised anomaly segmentation accuracy by allowing the model to be guided by expert input. Future work could explore training the model to be sensitive to specific types of anomalies while disregarding others, enabling selective segmentation. Such advancements would allow the model not only to detect anomalies but also to classify them to some extent, thereby adding greater value to clinical decision-making and facilitating more tailored medical interventions.
The proposed approach is particularly well-suited for addressing anomalies that are distinguishable based on intensity differences. Anomalies that present clear intensity variations compared to surrounding tissues are efficiently handled by this framework, resulting in high segmentation accuracy. However, the approach is less effective when dealing with anomalies that do not have distinctive intensity contrasts but instead differ in shape or structural characteristics. In such cases, the proposed model does not show evident superiority to differentiate between normal and abnormal shapes, as the generative process is primarily guided by intensity-based information. Future work could explore combining intensity-based anomaly detection with additional spatial or shape features to extend the model’s applicability to a broader range of complex anomalies.
Declaration of Competing Interest
The authors report no conflicts of interest.

### v2

@@
-Synomaly Noise and Multi-Stage Diffusion: A Novel Approach for Unsupervised Anomaly Detection in Ultrasound Imaging
+Synomaly Noise and Multi-Stage Diffusion: A Novel Approach for Unsupervised Anomaly Detection in Medical Images
 Abstract
-Ultrasound (US) imaging is widely used in routine clinical practice due to its advantages of being radiation-free, cost-effective, and portable.
-However, the low reproducibility and quality of US images, combined with the scarcity of expert-level annotation, make the training of fully supervised segmentation models challenging.
+Anomaly detection in medical imaging plays a crucial role in identifying pathological regions across various imaging modalities, such as brain MRI, liver CT, and carotid ultrasound (US).
+However, training fully supervised segmentation models is often hindered by the scarcity of expert annotations and the complexity of diverse anatomical structures.
 To address these issues, we propose a novel unsupervised anomaly detection framework based on a diffusion model that incorporates a synthetic anomaly (Synomaly) noise function and a multi-stage diffusion process.
@@
 Notably, the unsupervised anomaly detection model is trained purely on healthy images, eliminating the need for anomalous training samples and pixel-level annotations.
-We validate the proposed approach on carotid US, brain MRI, and liver CT datasets.
+We validate the proposed approach on brain MRI, liver CT datasets, and carotid US.
 The experimental results demonstrate that the proposed framework outperforms existing state-of-the-art unsupervised anomaly detection methods, achieving performance comparable to fully supervised segmentation models in the US dataset.
-Additionally, ablation studies underline the importance of hyperparameter selection for Synomaly noise and the effectiveness of the multi-stage diffusion process in enhancing model performance.
+Ablation studies further highlight the contributions of Synomaly noise and the multi-stage diffusion process in improving anomaly segmentation.
+These findings underscore the potential of our approach as a robust and annotation-efficient alternative for medical anomaly detection.
 Code: https://github.com/yuan-12138/Synomaly.
 keywords:
-Ultrasound Image Analysis, Unsupervised Anomaly Detection, Diffusion Model, Medical Image Analysis, Carotid imaging
+[…]
 1 Introduction
-Ultrasound (US) plays a crucial role in daily medical practice.
-It is commonly used for the screening and early assessment of internal organs and vasculature.
-Compared to computed tomography (CT) and magnetic resonance imaging (MRI), US is more frequently utilized for routine examinations.
-CT, although efficient in providing high-quality images, involves radiation exposure, which makes it less suitable for regular check-ups.
-MRI, while providing detailed soft tissue contrast, is often cost- and time-intensive for routine use.
-Consequently, physicians typically reserve CT and MRI for cases where a patient has already exhibited specific symptoms or conditions.
-As a result, CT and MRI are more commonly associated with the imaging of diseased cases.
-In contrast, US is radiation-free, cost-effective, and portable, making it an ideal modality for regular monitoring.
-Therefore, the US images collected are often from non-diseased individuals, making it as an ideal modality to learn the physiology of healthy tissues.
-However, the inherent challenges of low reproducibility and poor image quality in US imaging make it difficult to gather large amounts of diseased data for training of deep neural networks.
-This challenge is particularly evident in conditions like stroke, a leading cause of mortality worldwide, with ischaemic stroke accounting for about 65.3% of all cases [16].
-This type of stroke is caused by the blockage of a blood vessel supplying the brain, and carotid atherosclerosis is a common underlying cause [22].
-US is frequently used for the detection and monitoring of carotid plaques.
-One important indicator to assess the risk of stroke for atherosclerosis patients is total plaque volume [1].
-One pre-requisite for such measurement is an accurate segmentation of the carotid plaque in US images.
-However, unlike healthy anatomy, plaques vary significantly in shape and location, making their detection more challenging.
-This variability means that high-quality datasets are crucial for training segmentation models to reliably identify plaques.
-Without such datasets, models may struggle to generalize and correctly segment plaques in diverse cases.
-Because of ethical and privacy issues, it is not an easy practice to gather large amount of diseased US data.
-To tackle such challenge, unsupervised anomaly detection strategy appears to be an effective solution.
+Anomaly detection in medical imaging is a common task in medical image analysis, where pathological regions are identified for further evaluation and classification.
+For example, detecting suspicious tumor regions in the brain and liver from magnetic resonance imaging (MRI) or computed tomography (CT) scans, or identifying carotid plaques in ultrasound (US) images, is essential for assessing disease severity and guiding clinical decisions.
+Traditionally, these tasks have been performed manually by radiologists and sonographers—a process that is time-consuming, labor-intensive, and prone to inter-operator variability.
+With recent advances in deep learning, automated image analysis has emerged as a powerful tool to assist medical professionals in detecting and evaluating pathologies more efficiently [30, 51, 29, 11, 27, 32].
+However, the effectiveness and robustness of deep learning models depend heavily on training data quality.
+In the natural image domain, data collection is relatively straightforward, with images readily available online or captured using consumer-grade cameras.
+Annotations can be performed by large groups of non-experts, enabling rapid labeling of extensive datasets at a low cost.
+In contrast, medical imaging data is restricted by privacy regulations and ethical considerations, requiring institutional approval and de-identification [31].
+Annotation is costly and time-consuming, as it demands specialized expertise from physicians [12].
+Additionally, inter-operator variability in labeling can introduce bias, further limiting model generalization.
+To tackle this challenge, unsupervised anomaly detection approach appears to be an effective solution.
 Similar to how humans learn about diseased tissues, unsupervised anomaly detection method tends to distinguish the anomaly tissue from the normal anatomy by learning from healthy structures [21].
-Generative models are commonly utilized to generate synthetic counterfactual healthy samples based on diseased images.
-Since the model is purely trained on healthy data, the need for pixel-level labels is eliminated.
-Considering the high noise ratio and low contrast of US compared to CT and MRI, the labelling process for US sometimes introduces unavoidable ambiguities, thus complicating the fully-supervised training process [15].
+Generative models are commonly utilized to synthesize counterfactual healthy samples from diseased images, enabling anomaly detection without requiring pixel-level annotations [15].
+This eliminates the need for extensive manual labeling, which can be challenging due to variations in image quality, contrast, and noise across different imaging modalities.
+For example, the high noise ratio and low contrast of US make the labelling process for US sometimes introduces unavoidable ambiguities, thus complicating the fully-supervised training process [20].
 However, such drawback does not concern the unsupervised anomaly detection method.
 Furthermore, it also provides the possibility to detect unseen or rare anomaly cases, which are not well represented in the dataset for fully-supervised models.
-Nonetheless, unsupervised anomaly detection is less common in US due to the presence of speckle noise, shadows, and acoustic artifacts, which make it challenging to differentiate anomalies from healthy tissues.
-Such complexity of US images necessitates the development of more advanced generative models.
-Diffusion models have proven effective in modeling complex data distributions, resulting in higher image sample quality compared to generative adversarial networks (GANs) and variational autoencoders (VAEs) [12, 10].
+Despite its advantages, the effectiveness of unsupervised anomaly detection largely depends on the generative model’s ability to preserve healthy structures and accurately reconstruct pathological regions as their healthy counterparts.
+This underscores the need for further exploration of advanced generative models to enhance performance.
+Diffusion models have proven effective in modeling complex data distributions, resulting in higher image sample quality compared to generative adversarial networks (GANs) and variational autoencoders (VAEs) [17, 14, 43].
 These advantages are particularly valuable for unsupervised anomaly segmentation, especially in cases of small datasets, leading to growing interest in applying diffusion models to this application domain [35].
-Most recent works have also reported initial attempts of applying diffusion models for unsupervised anomaly detection in US [34, 32], specifically exploring their effectiveness for detecting anomalies in fetal brain US.
+Recent works have also reported several attempts of applying diffusion models for unsupervised anomaly detection in fetal brain US [47, 45], brain MRI [40], and brain CT [23].
 The performance of diffusion model is closely linked to the choice of corrupting noise, especially for medical image anomaly detection [34].
 Superior performance over naive pixel-wise Gaussian noise has been reported when applying Simplex noise [61] and Coarse noise [33] for anomaly detection of brain MRI.
-In order to develop a robust unsupervised anomaly detection model for US images, in this work, we propose Synomaly noise, which stands for synthetic anomaly noise based on statistical disease morphology prior.
+In order to develop a robust unsupervised anomaly detection model for medical images, in this work, we propose Synomaly noise, which stands for synthetic anomaly noise based on statistical disease morphology prior.
 Different from the aforementioned noise functions, Synomaly noise is specifically designed for unsupervised medical anomaly detection by corrupting healthy images with synthetic anomalies during training.
@@
 3.
-This is the first work to apply a diffusion model for unsupervised anomaly detection in carotid US images.
-We demonstrate the effectiveness of the proposed Synomaly noise and multi-stage diffusion process on a US carotid plaque dataset.
-To further validate the performance of the proposed approach, we conducted experiments on public brain MRI and liver CT datasets, achieving superior performance over state-of-the-art methods in detecting brain tumors and liver diseases, respectively.
+We conduct extensive validation across multiple medical imaging modalities, including brain MRI, liver CT, and carotid ultrasound (US) datasets.
+To the best of our knowledge, this is the first work to apply a diffusion model for unsupervised anomaly detection in carotid US images.
+Our approach achieves superior performance compared to state-of-the-art methods.
 The results of the three most common medical image modalities (US, CT, and MRI) demonstrate the potential of the proposed method in wide application.
-To facilitate the reproducibility of this work, the code will be released on this website upon acceptance.
+To facilitate the reproducibility of this work, the code will be released on this website upon acceptance.111Codes: https://github.com/yuan-12138/Synomaly
 2 Related Work
@@
 2.1 Unsupervised Anomaly Detection Using Autoencoders
-Autoencoders (AEs), since their introduction, have been widely used for unsupervised representation learning.
-The principles behind AE-based unsupervised anomaly detection/segmentation is to learn a latent space distribution that represents healthy anatomies, identifying abnormalities as outliers.
-Consequently, the reconstructed images preserve the healthy regions while altering the anomalous parts, making it easier to detect and segment the abnormalities.
-Different variants of AEs have been applied to different anomaly detection scenarios.
-[40] implemented a conditional VAE to perform anomaly segmentation in brain MRI, while [35] utilized a Bayesian convolutional AE to locate the abnormalities in CT.
-By combining VAE with a self-supervised image representation learning network, [29] proposed SSL-AnoVAE to tackle the anomaly staging and segmentation of retinal images.
-Deviated from the reconstruction training approaches, [44] adopted an inpainting strategy to train a context-encoding VAE for anomaly detection in brain MRI.
-In order to increase the ability of AEs in modeling complex high-dimensional data distributions, adversarial strategies have been introduced [21].
-[5] integrated GAN with VAE and proposed AnoVAEGAN to perform anomaly detection for brain MRI.
+Autoencoders (AEs) have been widely used for unsupervised anomaly detection by learning latent representations of healthy anatomy and identifying deviations as anomalies.
+The reconstructed outputs typically preserve normal regions while modifying anomalous ones, aiding detection and segmentation.
+Various AE variants have been applied across modalities [1, 41, 55].
+For example, [59] used a conditional VAE for brain MRI segmentation, and [48] applied a Bayesian convolutional AE to CT.
+[39] proposed SSL-AnoVAE by combining VAE with self-supervised learning for retinal images.
+Instead of standard reconstruction, [64] adopted a context-encoding VAE trained via inpainting.
+To improve AE capacity for modeling complex distributions, adversarial methods have been introduced.
+[5] combined VAE with GAN and proposed AnoVAEGAN for brain MRI anomaly detection.
 Apart from reconstruction-based approaches, restoration-based methods offer an alternative.
 Instead of directly passing input images through AE network to generate counterfactual reconstructions, restoration methods focus on finding the closest representation of the abnormal image in the learned latent space of normal images.
-These methods project the abnormal image onto the normal manifold and transform it into its closest healthy version.
+By projecting the abnormal image onto the normal manifold, the model transforms it into its most similar healthy counterpart [25].
 Anomalies are then detected by comparing the original input with the restored image.
@@
 The restoration process of anomaly samples is performed through an iterative gradient ascent method by Maximum-A-Posteriori estimation.
+2.2 Unsupervised Anomaly Detection Using Synthetic Anomalies
+Another mainstream direction in unsupervised anomaly detection is the synthetic anomaly approach, where the goal is to generate anomalous-looking images from normal ones to train a discriminative model.
+For example, CutPaste [37] creates pseudo-anomalies by cutting and pasting patches within a normal image to simulate realistic-looking defects.
+These synthesized images are then used to train a model to directly predict anomaly maps or scores.
+To improve realism, Natural Synthetic Anomalies (NSA) [54] employ Poisson image editing to seamlessly blend scaled patches from different images, enhancing the variability and natural appearance of the simulated anomalies.
+Similarly, in the medical domain, [58] applied this paradigm to chest X-ray and US images using Poisson image interpolation, generating anomalies that mimic real pathological structures.
+However, these methods do not produce counterfactual reconstructions and are dependent on the fidelity and realism of the generated pseudo-anomalies [57].
+If the true anomalies have highly out-of-distribution appearances, these approaches may struggle to generalize.
+Fundamentally, such methods attempt to model what anomalies might look like, rather than learning what healthy tissue should look like, as is done in reconstruction-based methods.
 2.3 Unsupervised Anomaly Detection Using Denoising Models
-Recent advances in diffusion models have demonstrated superior generative performance across various medical imaging modalities [13, 43, 14, 24].
-This enhanced performance has also proven beneficial for unsupervised anomaly detection in medical imaging, enabling more accurate identification of abnormalities [27, 28].
-[42] proposed AnoDDPM, where a Denoising Diffusion Probabilistic Model (DDPM) is utilized to reverse the diffusion process and generate counterfactual images.
-A partial diffusion process is introduced to better conserve the fine details of the input images and further increase the sample time efficiency during inference.
-They also replaced the common Gaussian noise with Simplex noise for image corruption and achieved better results when testing on brain MRI images.
-To avoid the accumulated errors during iterative denoising process and increase the sample efficiency of the diffusion process, [36] implemented latent diffusion model for anomaly detection in brain MRI and CT.
-The noising and denoising process is then executed in the latent space rather than directly on the images, which can significantly lower the dimensionality and improve efficiency.
-[41] utilized a Denoising Diffusion Implicit Model for medical anomaly detection, which enables a much faster sampling than DDPMs.
-The diffusion model is trained on both healthy and diseased images, and during inference, the generation of healthy counterparts for diseased images is guided by a pretrained classifier.
-Instead of providing direct guidance during the denoising process, [8] introduced THOR to provide implicit guidances through intermediate anomaly maps.
-This approach allows the model to focus on potential anomaly regions while preserving more details in the healthy areas, thereby reducing the false positive rate in anomaly detection.
-Similarly, [19] combined DDPM with ACAT [18], a saliency map that can approximately capture the pathological regions in the input image.
-At each denoising step, the regions identified by the ACAT saliency map are processed separately enabling targeted modification of potential diseased areas while preserving healthy regions.
-As a result, they demonstrated enhanced erasure performance of the anomalous regions in experiments on brain MRI and CT datasets.
-Considering the high computational burden of the diffusion process for high-resolution images like MRI, [6] proposed a patch-based diffusion model, which reconstructs input images patchwise, thereby reducing network complexity.
+Recent advances in diffusion models have shown strong generative performance across medical imaging modalities [18, 62, 19, 28], which has also benefited unsupervised anomaly detection by enabling more accurate identification of abnormalities [35, 38].
+[61] introduced AnoDDPM, a DDPM-based method for generating counterfactual images using a partial diffusion process to preserve fine image details and improve inference speed.
+They also adopted Simplex noise over Gaussian for better results on brain MRI.
+To avoid accumulated errors and improve sampling efficiency, [49] proposed a latent diffusion model, operating in compressed latent space rather than image space.
+[60] used a Denoising Diffusion Implicit Model, enabling faster inference.
+Their model was trained on both healthy and diseased images, with generation guided by a pretrained classifier.
+[9] proposed THOR, which incorporates anomaly maps during inference to focus on suspicious regions while preserving healthy structures.
+[23] combined DDPM with ACAT [22], using saliency maps to isolate and process potential anomalies separately, improving anomaly erasure in brain MRI and CT.
+To address the high computational cost for high-resolution images like MRI, [6] proposed a patch-based diffusion model that reconstructs images patchwise, reducing complexity.
 Given the nature of diffusion process, the corruption of input images will inevitably lead to the loss of details of the original images.
@@
 3 Method
-The performance of diffusion models in unsupervised anomaly detection is closely related to the design of the noising function.
-An effective noise function should efficiently corrupt the original image such that, after the noising process, the anomalous regions become indistinguishable from the normal ones.
-This allows the model to replace anomalies with normal tissues during the denoising process.
-However, the classic noise functions face challenges with larger or particularly distinct anomalies, such as very bright or dark regions.
-In such cases, many noise steps may be required to sufficiently blur the anomaly and make it indistinguishable from the surrounding healthy tissue.
-This extensive noise addition can lead to significant loss of structural image information, resulting in blurred and inaccurate image reconstruction.
+The performance of diffusion models in unsupervised anomaly detection is closely tied to the design of the noising function.
+An effective noise function should corrupt the image such that, after noising, anomalous regions become indistinguishable from normal tissue—allowing the model to replace them with healthy structures during denoising.
+However, classic noise functions struggle with large or highly distinct anomalies, e.g., very bright or dark regions, which often require many noise steps to obscure.
+This heavy corruption can degrade structural details and lead to blurred, inaccurate reconstructions.
 This introduces a precision-recall trade-off: Using fewer noise steps preserves structural integrity and achieves high reconstruction accuracy for healthy tissues, but it increases the risk of failing to detect distinct anomalies, leading to a higher rate of false negatives - characterized by high precision and low recall.
@@
 1(a).
-This allows the model to refine its denoising strategy specifically for reconstructing normal tissue from noised regions that resemble pathological anomalies.
-During inference, standard Gaussian noise is applied to the images, causing the abnormal regions to resemble the noised artificial anomalies encountered during training.
-This similarity enables the model to efficiently generate counterfactual healthy tissues and effectively replace anomalies with normal tissue during the denoising process.
+This enables the model to refine its denoising strategy for reconstructing normal tissue from corrupted regions resembling real corrupted anomalies.
+During inference, standard Gaussian noise is applied, making abnormal areas resemble the synthetic anomalies seen during training.
+This similarity allows the model to efficiently generate counterfactual healthy tissue and replace anomalies during denoising.
+Importantly, the goal of Synomaly noise is not to produce anatomically realistic lesions, but to create synthetic anomalies that mimic the visual appearance of real anomalies after noise corruption.
 To further improve the conservation of fine details during diffusion process, we employ an iterative anomaly erasure approach, as shown in Fig.
@@
 Instead of merely adding noise until anomalies blend into healthy tissues, Synomaly strategically integrates synthetic anomalies into the noise model during training.
-The core concept is to train the diffusion model to remove synthetic anomalies in healthy images that mimic real-world scenarios, thereby improving the model’s ability to detect and remove actual anomalies during inference.
-The key advantage of Synomaly in comparison to the other noise functions is its ability to utilize disease-specific information (morphology or appearance of the disease) during model training, which greatly enhances anomaly detection performance based on our experiments.
+The core idea is to train the diffusion model to remove synthetic anomalies embedded in healthy images that resemble real pathological patterns, thereby enhancing its ability to detect and remove true anomalies at inference.
 Fig.
@@
 Specifically, increasing the blurring variance () results in larger fragments of anomalies, while higher threshold values () reduce the number and size of anomalies by excluding lower Gaussian values.
-To further refine the simulation of real anomalies, an optional mask can be applied to constrain the anomaly shapes to relevant regions.
+To further refine the simulation of real anomalies, an optional mask can be applied to ensure that synthetic anomalies are placed within relevant anatomical regions rather than in background areas.
+This mask is particularly useful for imaging modalities where the foreground and background are not clearly separable or where anomalies are expected to occur within specific structures.
 Once the synthetic anomaly shapes are defined, synthetic anomalies are added to the Gaussian background noise at their respective locations.
@@
 Most of the existing works on diffusion models for anomaly detection typically suggest performing the diffusion process only once during inference, which is termed as “single-stage” approach.
-However, this approach is usually insufficient when dealing with anomalies that are unique in shape or intensity.
-Often, a single-stage cycle cannot fully remove anomalies, especially when many noise steps are required to achieve good anomaly detection.
-While increasing the number of noise steps can improve anomaly removal, it can also degrade structural information and eventually lead to blurred and inaccurately reconstructed images.
+However, this is often inadequate for anomalies with distinct appearance.
+A single pass may not fully remove such anomalies, especially when many noise steps are needed.
+While increasing noise steps can improve anomaly removal, it also degrades structural detail, resulting in blurred and inaccurate reconstructions.
 Therefore, we propose a “multi-stage diffusion process”.
@@
 3(a), starting with the original input image , the process begins by adding noise for a predefined number of noise steps, followed by denoising.
-Then, the denoised image is fused with the original image to have the image with reduced abnormality for the next stage.
-This fusion process is detailed in Fig.
-3(b), where the computed anomaly mask in the current stage is used to provide guidance.
-This process is repeated over multiple stages until the generated anomaly map converges.
-To prevent excessive iterations, the maximum number of iterations is set to five.
+Based on the denoised image , the anomaly mask is generated (Fig.
+3(b)) by thresholding the difference between blurred versions of the denoised and the original image.
+Using this mask, masked fusion (Fig.
+3(c)) combines anomalous regions from with the background from the original image , producing a new input for the next stage.
+This process iterates until the sum of detected anomaly pixels stabilizes, meaning that the relative change between consecutive anomaly masks falls below a threshold of 0.01.
+As a result, the number of diffusion stages at inference is not a manually set hyperparameter but is automatically determined based on the convergence of the anomaly mask.
+To prevent excessive computations, a maximum of five iterations is imposed, as further iterations typically yield negligible improvements.
 This approach offers several benefits.
-By using fewer noise steps at each stage, the structural integrity of the image is better preserved, leading to more accurate reconstructions.
+By decomposing a large noise step into multiple smaller steps across stages, the structural integrity of the image is better preserved, leading to more accurate reconstructions.
 Additionally, performing multiple inference stages with sequential denoising allows even distinct anomalies to be progressively removed and detected more effectively.
-Since different noise steps model different image features [23] and the goal of medical anomaly detection is to focus on finer details without excessively corrupting the entire image, fewer noise steps are particularly relevant for detecting subtle anomalous regions.
+Since different noise steps model different image features [26] and the goal of medical anomaly detection is to focus on finer details without excessively corrupting the entire image, fewer noise steps are particularly relevant for preserving details from non-anomalous regions.
 To further reduce the false positive rate in the generated anomaly mask, we propose a “masked fusion” strategy, as shown in Fig.
@@
 The detailed process is summarized in Algorithm 1.
-4 Experiments
+4 Experiments and Discussions
 In this section, we evaluate the performance of the proposed model using various datasets.
@@
 4.1 Datasets
-In this work, three datasets are used, namely, a brain MRI dataset, a liver CT dataset, and a carotid US dataset. The details of each dataset are provided below.
-4.1.1 Brain MRI: BraTS Dataset
-The BraTS2023 dataset originates from the international Brain Tumor Segmentation (BraTS) Challenge [2, 3, 31], which is a well-established benchmark in the field of medical imaging.
-It comprises 3D brain MRI scans from patients diagnosed with brain tumors, providing a rich and diverse set of data for developing and evaluating segmentation models.
-The pixel-wise annotations classify tissues into the four categories healthy tissue, GD-enhancing tumor (ET), peritumoral edematous/invaded tissue (ED) and necrotic tumor core (NCR).
-For the purposes of this work, all three tumor-related classes (ET, ED, and NCR) are grouped into a single “anomalous” category, forming two classes: healthy and anomalous tissue.
-Since the focus is on tumor detection in 2D images, only axial brain slices are considered in this work and the 2D tumor ground truth segmentation is extracted from the annotated 3D classification volumes.
-Given that ET, ED and NCR are most commonly found in the cerebral hemispheres, the brain MRI dataset is further refined to include only the central slices.
-For anomaly detection, the FLAIR sequence is used.
-All selected slices are resized to pixels to have unified dimensions.
-anomalous slices are randomly selected as the test set.
-In total, the dataset consists of healthy training samples and anomalous test samples.
-To help intuitive understanding, a few representative samples are visualized in Fig.
-4(b).
-4.1.2 Liver CT: LiTS Dataset
-The Liver Tumor Segmentation Benchmark (LiTS) dataset [9] includes abdominal CT volumes, each with detailed annotations of the liver and its tumors.
-The number of tumors per volume ranges from to , containing both primary and secondary tumors of diverse sizes and appearances.
-For the purposes of this work, all tumors are categorized as “anomalous” class.
-Each 3D abdominal CT volume contains varying axial slices, ranging from and .
-To prepare the data for model training and evaluation, the annotated segmentations are used to classify slices into healthy and anomalous cases.
-Further refinement is performed using size thresholding to isolate the relevant slices and all images are resized to pixels.
-In total, the CT dataset has healthy training images and anomalous test images.
-A set of representative samples is shown in Fig.
-4(c).
-These examples highlight the complexity of the anomaly detection task, as the abdominal region exhibits various appearances and the liver itself can differ in size, shape and location.
-The relatively small ratio between the anomaly and the chest volume can further cause class imbalance problems.
-These factors make it particularly challenging to accurately segment liver tumors, while preserving healthy tissue.
-This high diversity and complexity of the LiTS dataset make it a great benchmark for assessing the performance and generalizability of the proposed anomaly detection approaches.
-4.1.3 Carotid US Dataset
-The carotid artery US dataset includes US scans from healthy patients and patients with carotid plaques.
-The US images in the dataset are all captured from the cross-sectional view of the carotid artery.
-The acquisition was performed within the Institutional Review Board Approval by the Ethical Commission of the Technical University of Munich (reference number 244/19 S).
-Due to the nature of US imaging, B-mode appearance is affected significantly by contact point, force, angle, and US machine setting.
-So, to alleviate the effects caused by the changing image appearance, we crop the US image around the target anatomy, i.e., the carotid artery, to ensure the network only focuses on disease-cased image variations.
-In this study, we applied existing vessel tracking and segmentation methods to generate vessel masks, which are then used to crop the original full-view US images.
-Fig.
-4(a) presents the pipeline for vessel tracking and cropping procedure.
-To account for potential imperfections in vessel segmentation and to enhance model robustness, a small degree of randomness regarding the position and size of the vessel within the image is introduced in the cropping process, which also creates some variability in the training data.
-The ground truth annotations for plaques in the anomalous carotid artery samples are manually annotated under the supervision of our clinical experts specialized in vascular disease.
-All such annotations of vessels and anomalies were performed using the ImFusion Suite111https://www.imfusion.com/products/imfusion-suite.
-All US images are resized to pixels to standardize input dimensions.
-In total, the final dataset contains healthy training images and anomalous test images.
-The examples can be found in Fig.
-4(d).
+We evaluate the proposed method on three representative datasets spanning different modalities and anatomical regions: Brain MRI (BraTS 2023) [2, 3, 44], Liver CT (LiTS) [13], and Carotid US.
+The datasets comprise both healthy and anomalous cases, with Brain MRI containing 3743 healthy training slices, 1000 anomalous test slices, and 416 healthy test slices; Liver CT including 5820 healthy training images, 1000 anomalous test images, and 647 healthy test images; and Carotid US consisting of 7306 healthy training images, 545 anomalous test images, and 833 healthy test images.
+Additional details on data acquisition, preprocessing, and sample images are provided in Appendix A.
 | Method | Noise steps () | Kernel size () | Threshold () | ||||||
@@
 After the difference map is created by calculating the pixel-wise residuals between the blurred original image and blurred generated image , anomalies are identified by applying this anomaly threshold to the difference map, flagging regions with differences above this threshold as anomalies.
-In summary, the inference noise steps for partial diffusion (), the Gaussian kernel size for blurring (), and the anomaly threshold for anomaly selection () are the three hyperparameters that need to be carefully tuned for each dataset and diffusion noise function.
-To ensure a fair comparison between different models, the best inference hyperparameter is found for each model, representing the upper limit of performance.
+In summary, the inference noise steps for partial diffusion (), the Gaussian kernel size for blurring (), and the anomaly threshold for anomaly selection () are the three primary hyperparameters considered during inference.
+However, it is notable that the Gaussian kernel size and threshold value do not impact the quality of the generated counterfactual image.
+These parameters are used solely in the post-processing step to refine the anomaly mask by controlling the level of smoothing and the sensitivity of anomaly detection.
+In contrast, the number of noise steps plays a crucial role in determining the balance between preserving fine details and effectively removing anomalies during counterfactual image reconstruction.
+To ensure a fair comparison between different models, we determine the optimal inference hyperparameters for each approach, representing their best possible performance.
+As shown in Tab.
+1, the proposed method (DDPM+Synomaly+Multi-stage) shows minimal variation in noise steps across datasets, differing by only 100 steps, demonstrating its robustness to hyperparameter tuning.
 4.3 Training Details
-The performance of the proposed anomaly detection method is compared with six state-of-the-art (SOTA) methods: classic VAE, DDPM using Gaussian noise function [23], DDPM using Coarse noise function [26], DDPM using Pyramid noise function [20], DDPM using Simplex noise function [42], AutoDDPM [7], and DAE [25].
+The performance of the proposed anomaly detection method is compared with eight state-of-the-art (SOTA) methods: classic VAE [36], f-AnoGAN [52], NSA [54], DDPM using Gaussian noise function [26], DDPM using Coarse noise function [34], DDPM using Pyramid noise function [24], DDPM using Simplex noise function [61] (also known as AnoDDPM), AutoDDPM [8], and DAE [33].
 Fully supervised UNets [51] were also trained for each dataset as a reference.
 Half of the anomalous data of each dataset is selected as UNet training data.
-The VAE model and DAE model are trained for epochs using Adam optimizer with an exponentially decreasing learning rate from to by .
+| Method | Brain MRI BraTS23 | Liver CT LiTS | Carotid Plaque US | ||||||
+| DSC | Precision | Recall | DSC | Precision | Recall | DSC | Precision | Recall | |
+| VAE [51] | |||||||||
+| f-AnoGAN [52] | |||||||||
+| NSA [54] | |||||||||
+| DDPM+Gaussian [26] | |||||||||
+| DDPM+Corase [34] | |||||||||
+| DDPM+Pyramid [24] | |||||||||
+| DDPM+Simplex [61] | |||||||||
+| AutoDDPM [8] | |||||||||
+| DAE [33] | |||||||||
+| DDPM+Synomaly+Multi-stage | |||||||||
+| UNet [51]∗ | |||||||||
+Fully supervised method.
+The VAE model, f-AnoGAN, NSA model, and DAE model are trained for epochs using Adam optimizer with an exponentially decreasing learning rate from to by .
 All the DDPM models are implemented based on the popular guided diffusion approach by OpenAI [17].
@@
 All the models were trained using Nvidia A40 GPU.
-4.4 Results on Brain MRI
+During inference, DDIM [56] is applied to accelerate the inference process.
+4.4 Pixel-level Anomaly Detection Performance
+In this section, we evaluate the performance of the proposed method on pixel-level anomaly detection, comparing it with SOTA methods.
+The evaluation is conducted using anomalous images from the test set, where segmentation performance is measured to assess how effectively each method detects and localizes anomalies.
+Quantitative results are reported in Tab.
+2, while the visual results are shown in Fig.
+4.
+Since NSA is not a generative model and only outputs anomaly masks, its results are not shown in Fig.
+4.
+4.4.1 Results on Brain MRI
 For the anomaly detection in brain MRI images, the Synomaly noise function parameters are configured with a blurring variance of 3, a threshold value of 175, an anomaly direction of 1 and an intensity offset of 0.5.
 These Synomaly parameters were determined through a combination of known disease morphology and experimental validation.
+No mask is applied when generating Synomaly noise for brain MRI, as the brain occupies the majority of the image and is easily distinguishable from the background, reducing the risk of anomalies being misplaced in irrelevant areas.
 To ensure a fair comparison between methods, the optimal inference hyperparameters for each method were determined through a comprehensive grid search.
@@
 2.
-It is evident that the proposed Synomaly noise function consistently outperforms all other unsupervised methods across all metrics.
+It is evident that the proposed method consistently achieves the highest performance in terms of DSC and precision.
+While NSA yields the highest recall in the brain MRI dataset, its substantially lower precision suggests that it tends to over-segment, reducing its overall effectiveness.
 Fig.
@@
 However, changes in these areas can still be identified with a proper threshold.
-The VAE reconstruction is anomaly-free but appears blurry, losing important brain details.
+The reconstruction results of VAE and f-AnoGAN are anomaly-free but appears blurry, losing important brain details.
 Meanwhile, DDPMs trained with Gaussian noise and Pyramid noise fully remove the round anomaly in the top area of the image but fail to accurately reconstruct the lateral ventricle.
@@
 The proposed approach delivers the best anomaly segmentation mask compared to the other SOTAs.
-4.5 Results on Liver CT
+4.4.2 Results on Liver CT
 The configuration of Synomaly noise function for liver CT anomaly detection task is set as a blurring variance of 3, a threshold value of 175, an anomaly direction of -1 and an intensity offset of 1.
 Note that for this specific use case, the anomaly detection direction was set to -1 instead of 1, since the liver tumors appear darker than the surrounding healthy tissue, requiring an inversion of the detection parameter.
+The generated Synomaly noise is masked using the liver segmentation provided in the dataset, ensuring that synthetic anomalies are predominantly placed inside the liver region.
 The best inference hyperparameters are determined for each method by applying grid search.
@@
 2.
-In this scenario, the VAE completely fails to generate plausible counterfactual samples, as the complexity of CT images exceeds the generative capacity of the VAE model.
+In this scenario, the VAE, f-AnoGAN, and NSA completely fail to generate plausible counterfactual samples, as the complexity of CT images exceeds the generative capacity of the VAE model.
 As shown in Fig.
-5, the VAE-generated images are of low resolution with jagged borders.
+4, the generated images of VAE and f-AnoGAN are of low resolution with jagged borders.
 The DDPMs trained with Gaussian and Pyramid noise function can effectively erase the abnormalities but they also corrupt other parts of the image, resulting in a high false-positive detection rate.
@@
 The DDPM trained with the Simplex noise function demonstrated relatively good generation results, with liver tumors almost fully removed.
-The best performance was observed with Synomaly noise plus the multi-stage diffusion process, which achieved the highest scores across all metrics.
+The best performance was observed with Synomaly noise plus the multi-stage diffusion process, which achieved the highest scores in DSC and precision.
 The visual results also confirm its superior performance, as the anomalies are completely erased without any trace.
@@
 This can be attributed to the complexity of abdominal CT images, where the generative model must reconstruct not only the liver but also all other organs in the abdomen, making accurate segmentation considerably more challenging.
-4.6 Results on Carotid US
+4.4.3 Results on Carotid US
 The Synomaly noise function in this scenario was defined with parameters based on known disease morphology and experimental validation, which resulted in a blurring variance of 7, a threshold value of 150, an anomaly direction of 1 and an intensity offset of 0.5.
+A circular mask is applied to restrict synthetic anomalies within the vessel region, ensuring that the generated anomalies align with the expected anatomical distribution.
+The mask has a diameter of 90% of the image size and is centered within the image to approximate the vessel’s location.
 To identify the best inference parameters for each model, a comprehensive grid search was conducted across the parameters noise steps (), Gaussian kernel size () and anomaly threshold ().
@@
 Interestingly, the best performance apart from the proposed approach in this scenario is achieved by the VAE, with an average Dice score of .
-This can be attributed to the fact that, compared to brain MRI and liver CT, the structure of the carotid vessel is relatively simple, allowing the VAE to reconstruct a plausible plaque-free artery from a diseased one.
+This can be attributed to the fact that, compared to brain MRI and liver CT, the structure of the carotid vessel is relatively simple, allowing the generative models like GANs and VAEs to reconstruct a plausible plaque-free artery from a diseased one.
 Additionally, carotid plaques are more distinct and larger in size, which makes removing such anomalies with traditional SOTA noise functions challenging.
@@
 As observed in the fifth row of Fig.
-5, generated vessels often incorrectly interpret the inner border of the original diseased vessel as the outer border, leading to reconstructions of smaller-than-actual arteries.
+4, generated vessels often incorrectly interpret the inner border of the original diseased vessel as the outer border, leading to reconstructions of smaller-than-actual arteries.
 The DAE fails to effectively remove the plaques partly because the added noise during training does not sufficiently represent the anomalies.
@@
 As illustrated in the last subplot of the fifth row of Fig.
-5, the plaques are completely removed, and the artery’s borders are accurately reconstructed.
+4, the plaques are completely removed, and the artery’s borders are accurately reconstructed.
 Compared to the fully supervised method, the proposed unsupervised approach is only points lower in Dice score and even achieves a higher Recall.
@@
 Despite the existing domain gap, the proposed anomaly detection model still demonstrates comparable performance to the fully supervised method, highlighting its robustness and effectiveness in practice.
+We acknowledge that, to reduce complexity, a vessel tracking step was applied to localize the carotid artery region and minimize irrelevant background information.
+This preprocessing mitigates the high variability of US backgrounds, which can mislead generative models.
+Unlike CT and MRI, where background structures provide a stable anatomical reference, US backgrounds are inconsistent, making anomaly detection more challenging.
+Despite this simplification, the task remains difficult—even a fully supervised segmentation model trained on zoomed-in anomalous images achieved only 0.797 Dice, highlighting that anomaly detection remains a challenging task even after vessel localization.
+To demonstrate that vessel tracking can be achieved with minimal effort, we conducted an additional experiment using Segment Anything Model 2 (SAM2) [50] for automated carotid vessel tracking before applying the proposed anomaly detection method.
+Only the ground truth vessel mask of the first frame in each US sweep was provided to SAM2 for initialization, after which the model tracked the vessel across subsequent frames, and images were cropped based on the tracking results.
+Without any fine-tuning on US data, SAM2 successfully tracked the carotid region.
+Based on the tracking results, the final anomaly segmentation achieved a Dice score of .
+This suggests that modern large segmentation models can perform robust vessel tracking with minimal manual intervention.
+However, while vessel tracking benefits from strong shape priors, this may not generalize to more irregular structures like anomalies, where tracking remains challenging.
+These results highlight the feasibility of automated tracking as a practical preprocessing step for improving anomaly detection in US imaging.
+4.5 Image-level Anomaly Detection Performance
+In this section, we evaluate the performance of the proposed method on image-level anomaly detection, assessing its ability to distinguish anomalous images from healthy ones.
+The evaluation is conducted using both healthy and anomalous images in the test set, and performance is measured using the Area Under the Receiver Operating Characteristic Curve (AUROC).
+The results are compared across different imaging modalities to demonstrate the method’s effectiveness in various clinical scenarios.
+The AUROC scores achieved by the proposed method are 0.922 for brain MRI, 0.920 for liver CT, and 0.991 for carotid US, as illustrated in Fig.
+5.
+In the figure, the blue curve represents the AUROC for brain MRI, the orange curve corresponds to liver CT, and the green curve represents carotid US.
+These results indicate that the proposed method consistently performs well across different imaging modalities in distinguishing anomalous from healthy images.
+The particularly high AUROC score for carotid US (0.991) suggests that the method is highly effective in this modality, likely due to the more challenging appearance of anomalies compared to normal structures.
+While these results confirm the strong classification capability of the proposed method, pixel-level anomaly segmentation provides additional benefits by offering localized information about abnormal regions, which can improve interpretability and aid in detailed analysis.
+Additionally, pixel-level AD is inherently more challenging than image-level classification, as it requires the model to distinguish fine-grained abnormal structures from normal variations within an image.
+Nonetheless, the inclusion of image-level classification metrics further supports the overall effectiveness of the approach and provides a broader evaluation perspective.
 4.6 Ablation Study
-In this section, we conduct ablation studies on the hyperparameters of Synomaly noise and the multi-stage diffusion process.
-These experiments aim to provide readers with a clearer understanding of how these hyperparameters influence the final anomaly segmentation performance and highlight the importance of employing the multi-stage diffusion process.
-4.7.1 Hyperparameter Ablation for Synomaly Noise
-In this ablation study, we first examine the influence of the hyperparameters of Synomaly noise, as described in Sec.
-3.2.
-The performance of Synomaly noise is closely linked to its ability to represent the target anomalies.
-If the Synomaly noise generated during training resembles the real abnormalities encountered during inference, the trained model can more effectively erase the diseased areas during testing.
-The morphology of the generated Synomaly noise is controlled by two parameters: the blurring variance () and the threshold value ().
-To evaluate the influence of these parameters, several models were trained with different values of and on the carotid US dataset.
-The generated noise is normalized to the range [0, 255], and consequently, the threshold values are also within this range.
-First, we kept fixed at 150 and varied from 1 to 11.
-The results are presented in Fig.
-6(a), which shows that the best performance was achieved with a blurring variance of 7.
-Similarly, for , we conducted experiments by fixing at 7 while varying from 90 to 230.
-The results, as shown in Fig.
-6(b), indicate that the best performance was achieved when setting to 150.
-Some examples of the generated Synomaly noise are displayed below the plots.
-The value of controls the size of the synthetic anomaly regions, while determines the number of generated artificial anomaly areas.
-A smaller value results in smaller synthetic anomaly regions, whereas a lower value leads to a higher number of generated anomaly areas.
-However, in some cases, when choosing a smaller value, these generated regions sometimes overlap with each other, leading to fewer but larger anomalous areas.
-For the case of US plaque detection, the optimal choice is , , since this combination can best resemble the carotid plaque morphology based on our experiments.
+In this section, we conduct ablation studies to evaluate the impact of three key components of our method: the multi-stage diffusion process, Synomaly noise, and the hyperparameter selection for Synomaly noise.
+First, we analyze the effect of the multi-stage diffusion process by comparing it to the single-stage approach, assessing its role in improving anomaly removal while preserving fine details.
+Next, we evaluate the contribution of Synomaly noise to anomaly segmentation by comparing it with alternative noise functions.
+Finally, we examine the impact of hyperparameter selection for Synomaly noise, investigating how different configurations influence segmentation performance.
+These studies provide a deeper understanding of how each component contributes to the overall effectiveness of the proposed method.
 4.6.1 Ablation for Multi-stage Diffusion Process
@@
 In general, the number of noise steps at each stage for multi-stage inference is much smaller than the number of noise steps for single stage, which helps to preserve finer details.
-The quantitative results can be found in Tab.
-3 under the “Synomaly+Single-stage”, and the visual results can be found in Fig.
-7 on the right part.
+The quantitative ablation results can be found in Tab.
+3, and the visual results can be found in Fig.
+6.
+It is notable that the reconstructed images shown in Fig.
+6 are , used for anomaly mask generation.
+Due to the generative nature of diffusion, they may exhibit gradual intensity changes even in undetected anomaly regions.
+However, in the multi-stage framework, the input to the next stage is , where background regions are restored from the original image to prevent over-modification.
 | Metric | Brain MRI | Liver CT | Carotid US |
@@
 | DSC | |||
+| Precision | |||
 | Recall | |||
@@
 | DSC | |||
+| Precision | |||
 | Recall | |||
@@
 | DSC | |||
+| Precision | |||
 | Recall | |||
+| Brain MRI BraTS23 | Liver CT LiTS | Carotid Plaque US | |||||||
+| DSC | Precision | Recall | DSC | Precision | Recall | DSC | Precision | Recall | |
+| Stage 1 | |||||||||
+| Stage 2 | |||||||||
+| Stage 3 | |||||||||
+| Stage 4 | |||||||||
+| Stage 5 | |||||||||
 For the case of brain MRI dataset, the single-stage DDPM trained using Synomaly noise successfully erases the round anomaly but struggles with generating a realistic lateral ventricle, as shown in Fig.
-7(e).
+6(e).
 The results using the multi-stage diffusion process show a more realistic reconstruction, particularly for the lateral ventricle, compared to the single-stage process.
@@
 However, the comparison between Fig.
-7(a) and Fig.
-7(e) demonstrates that images generated using the multi-stage diffusion process are more realistic compared to those produced with a single diffusion step.
+6(a) and Fig.
+6(e) demonstrates that images generated using the multi-stage diffusion process are more realistic compared to those produced with a single diffusion step.
 For the case of liver CT, the multi-stage diffusion process has successfully increased the Dice by around 0.03 compared to the single-stage approach.
@@
 As depicted in Fig.
-7(f), the single-stage Synomaly method mostly removes the anomalies, but faint traces remain visible upon close inspection.
+6(f), the single-stage Synomaly method mostly removes the anomalies, but faint traces remain visible upon close inspection.
 In contrast, the multi-stage approach completely eliminates all the anomalies without a trace.
 Fig.
-6(b) shows how the abnormalities are progressively removed in each iteration of multi-stage diffusion process.
+7(b) shows how the abnormalities are progressively removed in each iteration of multi-stage diffusion process.
 For US carotid plaque anomaly segmentation, implementing the multi-stage diffusion process has improved the Dice score by approximately compared to the single-stage approach.
 The two visual examples in Fig.
-7(g) and Fig.
-7(h) also demonstrate that a single diffusion process cannot fully erase the anomalous regions.
+6(g) and Fig.
+6(h) also demonstrate that a single diffusion process cannot fully erase the anomalous regions.
 While the plaque in the upper part of the vessel is successfully removed, the larger plaque at the bottom of the vessel remains in the generated image.
@@
 Fig.
-7(c) shows the generated images at each stage, clearly illustrating how the plaques are iteratively removed from the original image.
+6(c) shows the generated images at each stage, clearly illustrating how the plaques are iteratively removed from the original image.
 Generally, distinct and large anomalies require more iterations to be completely erased.
 Fig.
-7(d) showcases another example involving five iterations, where the distinct plaque at the bottom of the vessel lumen is progressively removed over five iterations, while the smaller plaque at the top is mostly removed in four iterations.
+6(d) showcases another example involving five iterations, where the plaque at the bottom of the vessel lumen is progressively removed over five iterations, while the smaller plaque at the top is mostly removed in four iterations.
+The benefit of the multi-stage diffusion process is most evident in carotid US, where anomalies are more distinct and require stronger corruption to be removed.
+The multi-stage approach, combined with masked fusion, helps suppress these anomalies while preserving healthy regions.
+In contrast, for brain MRI and liver CT, where small noise levels are already sufficient to blur the anomaly, the improvement is less significant.
+To further evaluate the effectiveness of the proposed multi-stage diffusion process, we calculate the average DSC, precision, and recall of the intermediate anomaly masks at each stage, as summarized in Tab.
+4.
+For this analysis, we include only those test images that underwent five diffusion stages during inference.
+As shown in Tab.
+4, across the multi-stage diffusion process, recall consistently increases, while precision gradually decreases, and DSC improves and eventually stabilizes.
+This indicates that the model progressively captures more of the anomaly regions over successive stages, resulting in higher recall.
+The decrease in precision reflects a broader predicted mask that includes more false positives, but the gain in recall outweighs this, leading to improved DSC.
+By stage four and five, all metrics begin to converge, suggesting that the refinement process reaches a stable segmentation output.
 Additionally, we evaluated the performance without using the masked fusion technique, where the generated image from the previous iteration was directly used as input for the next iteration of the diffusion model.
@@
 By using the masked fusion technique, the original texture in non-diseased areas is preserved, thereby enhancing anomaly segmentation accuracy.
-5 Conclusion and Discussion
-In this paper, we presented a novel framework for unsupervised anomaly detection in US images using Synomaly noise and a multi-stage diffusion process.
-Synomaly noise was designed to add synthetic anomalies to healthy US images during training, enabling the model to effectively remove anomalies during inference.
+Lastly, we calculated the computational time for both single- and multi-stage diffusion processes.
+The proposed multi-stage diffusion process introduces additional computational steps compared to single-stage diffusion, resulting in an inference time of 67.65.7 milliseconds per image, compared to 13.85.6 milliseconds for the single-stage approach.
+The inference speed for other diffusion-based methods with single-stage diffusion is similar to our single-stage implementation.
+Notably, 67.6 milliseconds per image corresponds to approximately 15 frames per second, which is sufficient for most real-time applications.
+This ensures that the method remains computationally feasible for practical use while enabling effective anomaly detection.
+4.6.2 Ablation for Synomaly Noise
+To evaluate the effectiveness of Synomaly noise, we compare the performance of different noise functions when only a single-stage diffusion process is applied.
+The noise functions considered include Gaussian, Coarse, Pyramid, and Simplex noise, with the corresponding results presented in Tab.
+2 as DDPM+Gaussian, DDPM+Coarse, DDPM+Pyramid, and DDPM+Simplex, respectively.
+The performance of Synomaly noise in a single-stage diffusion process is reported in Tab.
+3 as Synomaly+Single-stage.
+For brain tumor segmentation, Synomaly noise achieved a Dice score of 0.738, outperforming other noise functions by at least 0.1.
+Similarly, for carotid plaque segmentation, Synomaly noise yielded a Dice score of 0.688, surpassing the best alternative by over 0.3.
+In liver tumor segmentation, Synomaly noise alone outperformed the best-performing SOTA noise function (Simplex noise) by 0.01 in Dice score, demonstrating its superiority across all datasets.
+While the performance gain in liver CT was less pronounced compared to the other datasets, Synomaly noise still consistently outperformed existing noise functions.
+These results highlight the advantages of Synomaly noise in generating realistic counterfactual images and enhancing anomaly segmentation performance.
+4.6.3 Hyperparameter Ablation for Synomaly Noise
+In this ablation study, we examine the influence of the hyperparameters of Synomaly noise, as described in Sec.
+3.2.
+The performance of Synomaly noise is closely linked to its ability to represent the target anomalies.
+If the Synomaly noise generated during training resembles the real abnormalities encountered during inference, the trained model can more effectively erase the diseased areas during testing.
+The morphology of the generated Synomaly noise is controlled by two parameters: the blurring variance () and the threshold value ().
+To evaluate the influence of these parameters, several models were trained with different values of and on the brain MRI, liver CT, and carotid US dataset, respectively.
+The generated noise is normalized to the range [0, 255], and consequently, the threshold values are also within this range.
+For brain MRI and liver CT, the threshold value () was fixed at 175, while the blurring variance () varied from 1 to 11.
+For carotid US, was fixed at 150, with also ranging from 1 to 11.
+The results are shown in Fig.
+7(a), where the red, green, and blue lines represent brain MRI, liver CT, and carotid US, respectively.
+It is evident that the optimal performance is achieved at for brain MRI and liver CT, whereas for carotid US, the best choice is .
+Similarly, for , we conducted experiments by fixing at 3 while varying from 90 to 230 for MRI and CT data.
+For US, was fixed at 7.
+The results, as shown in Fig.
+7(b), indicate that the best performance is achieved when setting to 175 for MRI and CT, while is the best choice for carotid US.
+Some examples of the generated Synomaly noise are displayed below the plots.
+The value of controls the size of the synthetic anomaly regions, while determines the number of generated artificial anomaly areas.
+A smaller value results in smaller synthetic anomaly regions, whereas a lower value leads to a higher number of generated anomaly areas.
+However, in some cases, when choosing a smaller value, these generated regions sometimes overlap with each other, leading to fewer but larger anomalous areas.
+According to this ablation study, the optimal parameters for brain MRI and liver CT data are , , while the best combination for carotid US is , , since these parameter choices can best resemble the morphologies of corresponding anomalies based on our experiments.
+To facilitate hyperparameter selection without an exhaustive grid search, we provide practical guidelines based on the relative size of anomalous regions.
+For a 128128 image, we recommend for small anomalies ( of the image), while for moderate anomalies (), such as brain and liver tumor a setting of is suggested.
+Intermediate cases () can use as a reasonable starting point.
+For larger anomalies (), such as carotid US plaques, is a suitable choice.
+These anchor points serve as a structured reference for hyperparameter selection.
+A smaller generates smaller synthetic anomaly regions, while a lower increases the number of generated anomaly areas.
+By following these guidelines, users can refine parameters accordingly, reducing computational overhead while maintaining robust anomaly detection performance.
+Due to the random size distribution of the generated synthetic anomalies (see Fig.
+7 lower half), a model trained with a single parameter setting is not strictly limited to detecting anomalies of a specific size but can still generalize to certain out-of-distribution cases.
+5 Conclusions
+In this paper, we presented a novel framework for unsupervised anomaly detection in medical images using Synomaly noise and a multi-stage diffusion process.
+Synomaly noise was designed to add synthetic anomalies to healthy images during training, enabling the model to effectively remove anomalies during inference.
 The multi-stage diffusion process was introduced to incrementally denoise the images, thereby preserving fine details and generating more realistic reconstructions.
@@
 Additionally, the ablation studies highlighted the importance of configuring reasonable hyperparameters for Synomaly noise, as its performance is closely tied to its ability to resemble the target anomalies.
-We also demonstrated the effectiveness of the multi-stage diffusion process and the masked fusion technique in improving segmentation quality, both of which contributed to enhanced model performance.
+Our experiments also showed that Synomaly noise alone improves anomaly segmentation performance compared to other noise functions, indicating its effectiveness in generating counterfactual images.
+We also evaluated the multi-stage diffusion process and masked fusion technique, which contributed to further improvements in segmentation quality.
+These findings suggest that both Synomaly noise and multi-stage diffusion play important roles in the overall performance of the proposed method.
 Overall, the proposed approach outperformed existing state-of-the-art methods for unsupervised anomaly detection, significantly narrowing the gap between unsupervised and fully supervised approaches.
@@
 Moreover, the proposed framework provides an opportunity for doctors to incorporate their prior knowledge of the disease’s appearance, thus enhancing unsupervised anomaly segmentation accuracy by allowing the model to be guided by expert input.
-Future work could explore training the model to be sensitive to specific types of anomalies while disregarding others, enabling selective segmentation.
+In the future, hyperparameter selection can also be automated for non-expert users by incorporating a few segmentation examples.
+Further research could also focus on could also explore training the model to selectively detect specific types of anomalies while disregarding others, enabling selective segmentation.
 Such advancements would allow the model not only to detect anomalies but also to classify them to some extent, thereby adding greater value to clinical decision-making and facilitating more tailored medical interventions.
@@
 In such cases, the proposed model does not show evident superiority to differentiate between normal and abnormal shapes, as the generative process is primarily guided by intensity-based information.
-Future work could explore combining intensity-based anomaly detection with additional spatial or shape features to extend the model’s applicability to a broader range of complex anomalies.
+Recent studies have explored integrating structural priors [10], synthetic perturbations [42], and language-guided representations [63] to enhance robustness in such scenarios.
+Future work could explore combining intensity-based anomaly detection with additional spatial or shape features to extend the model’s applicability to a broader range of complex anomalies [7].
 Declaration of Competing Interest
 The authors report no conflicts of interest.
+Statement
+During the preparation of this work the author(s) used ChatGPT in order to revise the manuscript for fluent and grammatically correct English.
+After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
+Appendix A. Dataset Details
+A.1 Brain MRI: BraTS Dataset
+The BraTS2023 dataset, from the international Brain Tumor Segmentation Challenge [2, 3, 44], is a widely used benchmark in medical imaging.
+It includes 3D brain MRI scans from tumor patients with pixel-wise annotations for healthy tissue, GD-enhancing tumor (ET), edematous/invaded tissue (ED), and necrotic tumor core (NCR).
+In this work, all tumor-related labels (ET, ED, NCR) are combined into a single “anomalous” class, yielding two categories: healthy and anomalous tissue.
+Since the focus is on tumor detection in 2D images, only axial brain slices are considered in this work and the 2D tumor ground truth segmentation is extracted from the annotated 3D classification volumes.
+Given that ET, ED and NCR are most commonly found in the cerebral hemispheres, the brain MRI dataset is further refined to include only the central slices.
+For anomaly detection, the FLAIR sequence is used.
+All selected slices are resized to pixels to have unified dimensions.
+anomalous slices are randomly selected as the test set.
+In total, the dataset consists of healthy training samples, anomalous test samples, and healthy test samples.
+For normalization, we first clip the intensity values at the 99th percentile to mitigate the effect of extreme outliers, and then apply min-max normalization.
+To help intuitive understanding, a few representative samples are visualized in Fig.
+A1(b).
+A.2 Liver CT: LiTS Dataset
+The Liver Tumor Segmentation Benchmark (LiTS) dataset [13] includes abdominal CT volumes, each with annotations of the liver and tumors.
+For the purposes of this work, all tumors are categorized as “anomalous” class.
+Each 3D abdominal CT volume contains varying axial slices, ranging from and .
+To prepare the data for model training and evaluation, the annotated segmentations are used to classify slices into healthy and anomalous cases.
+Further refinement is performed using size thresholding to isolate the relevant slices and all images are resized to pixels.
+In total, the CT dataset has healthy training images, anomalous test images, and healthy test images.
+To normalize the images, we first clip the intensity values at the 99th percentile to reduce the impact of extreme outliers, followed by min-max scaling to the [0,1] range.
+A set of representative samples is shown in Fig.
+A1(c).
+These examples highlight the complexity of the anomaly detection task, as the abdominal region exhibits various appearances and the liver itself can differ in size, shape and location.
+The relatively small ratio between the anomaly and the chest volume can further cause class imbalance problems.
+These factors make it particularly challenging to accurately segment liver tumors, while preserving healthy tissue.
+This high diversity and complexity of the LiTS dataset make it a great benchmark for assessing the performance and generalizability of the proposed anomaly detection approaches.
+A.3 Carotid US Dataset
+The carotid artery US dataset includes US scans from healthy patients and patients with carotid plaques.
+The US images in the dataset are all captured from the cross-sectional view of the carotid artery.
+The acquisition was performed within the Institutional Review Board Approval by the Ethical Commission of the Technical University of Munich (reference number 2025-60-S-CB).
+Due to the nature of US imaging, B-mode appearance is affected significantly by contact point, force, angle, and US machine setting.
+So, to alleviate the effects caused by the changing image appearance, we crop the US image around the target anatomy, i.e., the carotid artery, to ensure the network only focuses on disease-cased image variations.
+In this study, we applied existing vessel tracking and segmentation methods to generate vessel masks, which are then used to crop the original full-view US images.
+Fig.
+A1(a) presents the pipeline for vessel tracking and cropping procedure.
+To account for potential imperfections in vessel segmentation and to enhance model robustness, a small degree of randomness regarding the position and size of the vessel within the image is introduced in the cropping process, which also creates some variability in the training data.
+The ground truth annotations for plaques in the anomalous carotid artery samples are manually annotated under the supervision of our clinical experts specialized in vascular disease.
+All such annotations of vessels and anomalies were performed using the ImFusion Suite222https://www.imfusion.com/products/imfusion-suite.
+All US images are resized to pixels to standardize input dimensions.
+A min-max normalization is applied to normalize the images to [0,1].
+In total, the final dataset contains healthy training images, anomalous test images, and healthy test images.
+The examples can be found in Fig.
+A1(d).
+Appendix B. Additional Visual Results
+Fig. B1, B2, and B3 present additional visual results on some challenging cases in the three datasets.
+Fig.
+B1 shows three extra visual results from the brain MRI dataset.
+The first two examples illustrate failure cases of the proposed method.
+In the first, the model incorrectly identifies a small bright region in the upper part of the image as an anomaly.
+Increasing the value and decreasing the value may improve robustness by making the model less sensitive to small artifacts and more focused on larger anomalous regions.
+In the second case, the failure is primarily due to the out-of-distribution appearance of the input, which deviates significantly from the training distribution.
+Incorporating a broader variety of training samples may help mitigate such issues.
+Nevertheless, the proposed method still demonstrates superior performance compared to other SOTA approaches, even in these challenging cases.
+Fig.
+B2 presents three additional examples from the liver CT dataset.
+In the first example, the model fails to remove three lesions, likely mistaking them for small hepatic ducts, resulting in missed anomaly detection.
+Decreasing and increasing during training could improve sensitivity to such subtle lesions.
+In the second case, the gastrointestinal duct appears as an irregular structure in the lower left but is correctly reconstructed and not mistaken as an anomaly.
+However, a dark region in the liver is wrongly identified, causing false positives.
+The third example shows good localization, with most anomalies accurately detected.
+Despite these challenges, the proposed method consistently outperforms other SOTA approaches.
+Fig.
+B3 presents three additional examples from the carotid US dataset.
+In the first image, an artifact within the carotid lumen is incorrectly identified as an anomaly and removed in the generated counterfactual image, resulting in a false positive prediction.
+In the second case, the plaque exhibits a highly distinct boundary and appearance.
+While the proposed method successfully detects some smaller plaques in the upper region, the larger and more prominent plaque at the bottom is not fully removed.
+A similar issue is observed in the third example, where the distinct appearance of the plaque hinders its complete removal.
+To improve the model’s performance on such distinct anomalies, increasing the intensity offset during training can produce more distinct Synomaly noise, enhancing the model’s ability to detect and remove pronounced anomalies.
+Despite these challenges, the proposed method still demonstrates robust performance compared to other SOTA approaches.